### v1

# On the cardinality of measures of maximal relative entropy for smooth skew products

Abstract.
Let $\Omega$ and $M$ be compact smooth manifolds and let $\Theta:\Omega\times M\to\Omega\times M$ be a $\mathcal{C}^{1+\alpha}$ skew-product diffeomorphism over an Axiom A base. We show that $\Theta$ has at most countably many ergodic hyperbolic measures of maximal relative entropy. When $\dim M=2$, if $\Theta$ has positive relative topological entropy, then $\Theta$ has at most countably many ergodic measures of maximal relative entropy.
Key words and phrases:
Measures of maximal relative entropy; skew products; random dynamical systems; hyperbolic dynamical systems; thermodynamic formalism
2020 Mathematics Subject Classification:
28D20; 37D25; 37D35; 37H05
1. Introduction and Main results
For dynamical systems related by a factor map, the natural analogue of maximal entropy is the concept of maximal relative entropy. Let $(X,f)$ and $(Y,g)$ be continuous maps on compact metric spaces and let $\pi\colon X\to Y$ satisfy $g\circ\pi=\pi\circ f$. Write $\mathcal{M}(f)$ and $\mathcal{M}(g)$ for the sets of invariant Borel probability measures of $f$ and $g$, respectively. Given $\nu\in\mathcal{M}(g)$, a measure $\mu\in\mathcal{M}(f)$ is a measure of maximal $\nu$-relative entropy if $\pi_{*}\mu=\mu\circ\pi^{-1}=\nu$ and
$h_{\mu}(f)=\sup\bigl\{\,h_{\eta}(f):\eta\in\mathcal{M}(f),\ \pi_{*}\eta=\nu\,\bigr\}.$
Equivalently, if $h_{\nu}(g)<\infty$ the Abramov–Rokhlin formula states that $\mu$ maximises the conditional entropy
$h_{\eta}(f\mid\nu):=h_{\eta}\left(f\mid\pi^{-1}\mathcal{B}_{Y}\right)=h_{\eta}(f)-h_{\nu}(g)$
among all $\pi$-lifts $\eta$ of $\nu$. This relative viewpoint is captured by the relativised variational principle of Ledrappier–Walters and in Walters’ compensation function, extending the notion of relative entropy to relative equilibrium states [29, 47].
The general picture of measures of maximal relative entropy and relative equilibrium states is well advanced in symbolic dynamics, particularly for factor maps between shifts where $X$ is an irreducible subshift of finite type. Petersen–Quas–Shin proved that for each ergodic base measure $\nu$ on $Y$ there are only finitely many ergodic measures of maximal relative entropy on $X$, and they established explicit bounds on this number [39]. These bounds were improved in [2], introducing the notion of class degree. A subsequent work extended these advances to relative equilibrium states in [1]. A structure theorem for infinite-to-one factor codes $\pi\colon X\to Y$ between irreducible subshifts of finite type was proved in [48]: that there exists a sofic shift $Y_{1}$, such that $\pi=\pi_{2}\circ\pi_{1}$. This allowed Yoo to show that if $\nu$ is an equilibrium state on $Y$ for a sufficiently regular potential, then there is a unique lift on $X$ that is a relative equilibrium state [48].
Measures of relative maximal entropy and relative equilibrium states naturally arise in random dynamical systems. There, they are usually called quenched relative equilibrium states. Consider a probability-preserving base $(\Omega,\theta,\mathbb{P})$ and the skew-product map
$\Theta:(\omega,x)\in\Omega\times M\mapsto(\theta(\omega),\,T_{\omega}(x))\in\Omega\times M.$
Observe that when taking $X=\Omega\times M$, $Y=\Omega$, $f=\Theta,$ $g=\theta$, and $\pi(\omega,x)=\mathrm{proj}_{\Omega}(\omega,x)=\omega$ we retrieve the setting of the first paragraph under the assumption that $h_{\mathbb{P}}(\theta)<\infty.$
Foundational results of Kifer established existence and uniqueness of relative equilibrium states for random uniformly expanding transformations and identified quenched relative pressure with the principal Lyapunov exponent of the random transfer-operator cocycle [27]. Bogenschütz developed parallel notions of metric entropy and pressure for random dynamical systems and proved a variational principle, in the framework commonly used nowadays [11]. Thermodynamic formalism for random systems has since been developed across a variety of settings: random symbolic dynamics [12, 21, 19, 44, 43]; smooth expanding maps [27, 33]; non-uniformly smooth expanding maps [4, 45]; and non-uniformly hyperbolic random interval maps (closed and open, with discontinuities) [7, 5, 6]; see also the references therein.
Although the quenched relative thermodynamic formalism theory for random dynamical systems is extensive, results for systems generated by diffeomorphisms remain scarce outside the SRB setting (for SRB results, see [28, 10, 20, 3, 32] and the references therein). In particular, except in settings where Gibbs-type bounds are available, there is no general method for bounding the number of measures of relative maximal entropy, let alone the number of relative equilibrium states, in these settings. In this paper, we study how many measures of relative maximal entropy a skew product $\Theta(\omega,x)=(\theta(\omega),T_{\omega}(x))$ over an Axiom A base may admit. This question is part of an ongoing program to quantify the number of measures of maximal entropy across classes of systems.
In the setting of smooth diffeomorphisms, Newhouse proved that $\mathcal{C}^{\infty}$ diffeomorphisms on compact smooth manifolds always admit a measure of maximal entropy [34]. Buzzi later conjectured that for surface diffeomorphisms with positive topological entropy, there are at most countably many ergodic measures of maximal entropy [15, Conjecture 1]. Sarig’s groundbreaking symbolic coding for $C^{1+\alpha}$ surface diffeomorphisms confirmed this countability and also settled a conjecture of Katok on periodic-orbit growth [40]. This result was recently sharpened by Buzzi–Crovisier–Sarig, proving that $\mathcal{C}^{\infty}$ surface diffeomorphisms with positive entropy have only finitely many ergodic measures of maximal entropy, and exactly one in the transitive case, [17], which answered another of Buzzi’s conjectures [15, Conjecture 2].
In higher dimensions, Ben Ovadia extended Sarig’s coding to non-uniformly hyperbolic diffeomorphisms on arbitrary manifolds, yielding countability results for hyperbolic measures of maximal entropy [9]. More recently, Buzzi–Crovisier–Sarig established the strong positive recurrence (SPR) property, and via SPR obtained strong statistical properties for the associated invariant measures, including exponential decay of correlations and an almost-sure invariance principle [18]. Related results hold for certain two-dimensional fibre skew products with hyperbolic base dynamics, due to Marín–Poletti–Veiga, and for flows [31, 16, 30].
Our contribution is a relative analogue of these countable cardinality results. For $\mathcal{C}^{1+\alpha}$ skew–product diffeomorphisms $\Theta(\omega,x)=(\theta\omega,T_{\omega}(x))$ over an Axiom A base $(\Omega,\theta)$ and for any ergodic base-invariant measure $\mathbb{P}$, we prove that there are only countably many ergodic hyperbolic $\Theta$-invariant measures of maximal entropy relative to $\mathbb{P}$. In particular, when the fibre $M$ is a surface and $h_{\mathrm{top}}(\Theta\mid\nu)>0$, the set of ergodic $\mathbb{P}$-relative measures of maximal entropy is countable. The proof refines the symbolic coding methods of [40, 18] to the skew-product setting and combines them with the finiteness techniques for factor maps due to Petersen–Quas–Shin [39]. We expect the coding-and-factor approach developed here to be useful in random dynamics, and to provide a natural quenched perspective and new tools for constructing and counting relative equilibrium states.
Structure of the paper.
The paper is organised as follows. The remainder of Section 1 states the main results and presents an application. Section 2 recalls background from Pesin theory and reviews homoclinic classes. In Section 3 we construct a natural symbolic coding for the skew product $\Theta$ that preserves the base dynamics. Section 4 generalises the results of [39] to countable-state Markov shifts. Finally, Section 5 contains the proofs of the main results, Theorem A and Corollary B build upon the results established in Sections 3 and 4.
1.1. Measures of maximal relative entropy
Throughout this paper, $\Omega$ and $M$ will always denote compact smooth (boundaryless) manifolds. Let $\Theta$ be a $\mathcal{C}^{1+\alpha}$, $\alpha>0$, skew-product diffeomorphism
$\Theta:\Omega\times M$ $\to\Omega\times M$ (1)
$(\omega,x)$ $\mapsto(\theta\omega,T_{\omega}(x))$
where $\theta:\Omega\to\Omega$, $T_{\omega}:M\to M$ for each $\omega\in\Omega$. Moreover, we define
$\mathrm{proj}_{\Omega}:(\omega,x)\in\Omega\times M\mapsto\omega\in\Omega.$
In this paper, all measures are implicitly assumed to be Borel measures.
This paper concerns the amount of $\mathbb{P}$-relative entropy that $\Theta$ may admit. Below, we define the metric entropy via the Brin–Katok formula [14] since it is the shortest way to define it (see also [46, Chapter 4]). We also define the concept of $\mathbb{P}$-relative entropy for skew products via an extension of the Brin–Katok formula (see [50]).
Definition 1.1 (Metric entropy).
Let $(X,\mathrm{dist})$ be a compact metric space, $T:X\to X$ continuous, and $\mu$ a $T$-invariant probability measure. For $x\in X$, $\varepsilon>0$, $n\in\mathbb{N}$, define the Bowen ball
$B_{n}(x,\varepsilon):=\left\{y\in X;\ \mathrm{dist}(T^{k}(y),T^{k}(x))<\varepsilon\text{ for }0\leq k<n\right\}.$
The metric entropy of $\mu$ with respect to $T$ is
$h_{\mu}(T):=-\int_{X}\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{1}{n}\log\mu(B_{n}(x,\varepsilon))\ \mathrm{d}\mu=-\int_{X}\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}\log\mu(B_{n}(x,\varepsilon))\ d\mu.$
Definition 1.2 (Relative metric entropy).
Let $\Theta$ be as (1) and $\mu$ be a $\Theta$-invariant probability with $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$, and disintegrate $\mu(\mathrm{d}\omega,\mathrm{d}x)=\mu_{\omega}(\mathrm{d}x)\,\mathbb{P}(\mathrm{d}\omega).$ For $\omega\in\Omega$, $x\in M$, $\varepsilon>0$, $n\in\mathbb{N}$, write
$B_{n}^{\omega}(x,\varepsilon):=\left\{y\in M;\ \mathrm{dist}\left(T_{\omega}^{k}(y),T_{\omega}^{k}(x)\right)<\varepsilon\text{ for }0\leq k<n\right\},$
where $T_{\omega}^{k}(x):=T_{\theta^{k}\omega}\circ T_{\theta^{k-1}\omega}\circ\cdots\circ T_{\omega}(x).$ The $\mathbb{P}$-relative metric entropy of $\mu$ with respect to $\Theta$ is
$h_{\mu}(\Theta\mid\mathbb{P})$ $:=-\int_{\Omega}\int_{M}\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{1}{n}\log(\mu_{\omega}(B_{n}^{\omega}(x,\varepsilon)))\,\mu_{\omega}(\mathrm{d}x)\,\mathbb{P}(\mathrm{d}\omega)$
$=-\int_{\Omega}\int_{M}\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}\log(\mu_{\omega}(B_{n}^{\omega}(x,\varepsilon)))\,\mu_{\omega}(\mathrm{d}x)\,\mathbb{P}(\mathrm{d}\omega).$
We recall that if $h_{\mathbb{P}}(\theta)<\infty$, which will always be the case in this paper, we have that $h_{\mu}(\Theta\mid\mathbb{P})=h_{\mu}(\Theta)-h_{\mathbb{P}}(\theta)$ (see [24, Chapter 5, below equation 1.1.5]).
Definition 1.3 (Measures of maximal $\mathbb{P}$-relative entropy).
Let $\Theta$ be as (1) and fix a $\theta$-invariant probability measure $\mathbb{P}$. We define the $\mathbb{P}$-relative topological entropy as
$h_{\mathrm{top}}(\Theta\mid\mathbb{P}):=\sup\{h_{\nu}(\Theta\mid\mathbb{P});\ (\mathrm{proj}_{\Omega})_{*}\nu=\mathbb{P}\ \text{and}\ \nu\ \text{is a }\Theta\text{-invariant probability measure}\}.$
We say that a $\Theta$-invariant measure $\mu$ is a measure of maximal $\mathbb{P}$-relative entropy if
$h_{\mu}(\Theta\mid\mathbb{P})=h_{\mathrm{top}}(\Theta\mid\mathbb{P}).$
To formulate our results, we impose additional assumptions on the skew-product base $\theta:\Omega\to\Omega$ and the measure $\mathbb{P}$, namely, we assume $\theta$ is an Axiom A diffeomorphism and $\mathbb{P}$ is an ergodic $\theta$-invariant probability measure.
Definition 1.4.
Let $\Omega$ be a compact smooth manifold and $\theta:\Omega\to\Omega$ be $\mathcal{C}^{1}$ diffeomorphism. We say that $\theta$ is an Axiom A transformation if
  • (i)
    $\overline{\{p\in\Omega;\ \text{$p$ is a $\theta$-periodic point}\}}=NW(\theta)$, where
    $NW(\theta):=\left\{\omega\in\Omega\,\middle|\,\begin{aligned} &\text{for each open set }U\subset\Omega\text{ containing }\omega,\\ &\#\{n\in\mathbb{N};\theta^{n}(U)\cap U\neq\emptyset\}=\#\{n\in\mathbb{N};\ \theta^{-n}(U)\cap U\neq\emptyset\}=\infty\end{aligned}\right\};$
    and
  • (ii)
    the non-wandering set $NW(\theta)$ is a uniformly hyperbolic set, i.e. there exists $\chi>0$ and a direct sum $T_{\omega}\Omega=E^{s}(\omega)\oplus E^{u}(\omega)$ for all $\omega\in NW(\theta),$ and a uniform number $K>0$ such that for any $k\geq 0,$ and $\omega^{\prime}\in NW(\theta)$,
    $\mathrm{max}\left(\left\|\left.\mathrm{D}\theta^{k}\right|_{E^{s}(\omega^{\prime})}\right\|,\left\|\left.\mathrm{D}\theta^{-k}\right|_{E^{s}(\omega^{\prime})}\right\|\right)\leq Ke^{-\chi k}.$
1.2. Main results
Below, we state our main results.
Theorem A.
Let $\Omega$ and $M$ be compact smooth manifolds, and
$\Theta:\Omega\times M$ $\to\Omega\times M$
$(\omega,x)$ $\mapsto(\theta\omega,T_{\omega}(x))$
be a $\mathcal{C}^{1+\alpha}$ diffeomorphism, for some $\alpha>0$. Assume that $\theta:\Omega\to\Omega$ is an Axiom $A$ diffeomorphism and $\mathbb{P}$ an ergodic probability measure. Then, $\Theta$ admits at most countably many ergodic hyperbolic measures of maximal $\mathbb{P}$-relative entropy.
Theorem A is proved in Section 5.
We remark that Theorem A is a generalisation of the classical non-skew-product setting. Let $T:M\to M$ be a $\mathcal{C}^{1+\alpha}$ map, $\theta:\Omega\to\Omega$ be any smooth Axiom $A$ map, and consider the product map $\Theta(\omega,x)=(\theta(\omega),T(x))$. Fix a base point $\omega_{p}\in\Omega$ with $\theta(\omega_{p})=\omega_{p}$ and let $\mathbb{P}=\delta_{\omega_{p}}$. Then any $\Theta$-invariant measure $\mu$ satisfying $(\mathrm{proj}_{\Omega})_{*}\mu=\delta_{\omega_{p}}$ has the form $\mu=\delta_{\omega_{p}}\otimes\mu^{\prime}$, where $\mu^{\prime}$ is $T$-invariant on $M$, and
$h_{\mu}\big(\Theta\,\big|\,\delta_{\omega_{p}}\big)=h_{\mu^{\prime}}(T).$
Hence, measures of maximal $\mathbb{P}$-relative entropy for $\Theta$ correspond exactly to measures of maximal entropy for $T$. In particular, Theorem A yields the usual countability statement for ergodic hyperbolic measures of maximal entropy in the non–skew-product setting in conformance with [9, Theorem 0.5].
We stress that the hyperbolicity assumption in Theorem A is essential. Observe that the extension of $\theta$ by the identity map $\Theta(\omega,x)=(\theta(\omega),x)$ admits uncountably many ergodic measures of maximal $\mathbb{P}$-relative entropy; in this case, all such measures have zero $\mathbb{P}$-relative entropy. A similar example with positive $\mathbb{P}$-relative entropy is $\Theta:\Omega\times\mathbb{T}^{3}\to\Omega\times\mathbb{T}^{3}$ with
$\Theta(\omega,(x_{1},x_{2},y))=(\theta\omega,(2x_{1}+x_{2},x_{1}+x_{2},y)),$
which admits uncountably many ergodic measures of maximal $\mathbb{P}$-relative entropy.
Moreover, for fibre maps in dimension $n\geq 3$, the countability conclusion for hyperbolic measures of maximal entropy in Theorem A is sharp, i.e. there exists a $\mathcal{C}^{\infty}$ skew-product diffeomorphism $\Theta:\Omega\times\mathbb{T}^{n}\to\Omega\times\mathbb{T}^{n}$ satisfying the hypotheses of Theorem A which admits infinitely many ergodic hyperbolic measures of $\mathbb{P}$-relative maximal entropy. As before, take $\mathbb{P}=\delta_{p}$, where $p\in\Omega$ is a fixed point of the base dynamics and set $\Theta(\omega,x)=(\theta(\omega),T(x))$, where $T:\mathbb{T}^{n}\to\mathbb{T}^{n}$ admits infinitely many hyperbolic measures of maximal entropy. One possible construction of $T$ is as follows.
Define
$T:(x,s)\in\mathbb{T}^{n-1}\times\mathbb{S}^{1}\mapsto(A(x),F(s))\in\mathbb{T}^{n-1}\times\mathbb{S}^{1},$
where $A\in\mathrm{SL}(n-1,\mathbb{Z})$ is an Anosov toral automorphism, and $F\colon\mathbb{S}^{1}\to\mathbb{S}^{1}$ is the function
$F(s)=s+\delta\,e^{-\left(\frac{1}{s^{2}}+\frac{1}{(1-s)^{2}}\right)}\sin\left(\frac{2\pi}{s}\right)\pmod{1}$,
where $\delta>0$ is taken sufficiently small so that $F$ is a diffeomorphism and $F^{\prime}(s)>0$ for every $s\in\mathbb{S}^{1}$. One readily verifies that $F\in\mathcal{C}^{\infty}$ and
$F\left(\frac{1}{j}\right)=\frac{1}{j}\ \text{and }0<F^{\prime}\left(\frac{1}{j}\right)=1-2\delta\pi j^{2}\,e^{-\left(1+\frac{1}{(j-1)^{2}}\right)j^{2}}<1\ \text{for every }j\in\mathbb{N}_{>1}.$
Since circle diffeomorphisms satisfy $h_{\mathrm{top}}(F)=0$, we obtain that $h_{\mathrm{top}}(T)=h_{\mathrm{top}}(A)$. If $m$ is the (unique) measure of maximal entropy for $A$ (in fact, the Lebesgue measure on $\mathbb{T}^{n-1}$), then for each $j\in\mathbb{N}_{>1}$ the product $m_{j}\coloneqq m\otimes\delta_{1/j}$ is ergodic, hyperbolic, and satisfies $h_{m_{j}}(T)=h_{m}(A)=h_{\mathrm{top}}(T)$. Thus $T$ admits infinitely many ergodic hyperbolic measures of maximal entropy.
In dimension two, Theorem A can be strengthened.
Corollary B.
Under the assumptions of Theorem A if we assume that $M$ is a compact smooth surface and $h_{\mathrm{top}}(\Theta\mid\mathbb{P})>0$ then $\Theta$ admits at most countably many measures of maximal $\mathbb{P}$-relative entropy.
Corollary B is proved in Section 5.
In dimension two, it may be possible to strengthen the conclusion of “at most countably many measures of maximal $\mathbb{P}$-relative entropy” of Corollary B to “at most finitely many measures of maximal $\mathbb{P}$-relative entropy”, in analogy with the classical non-skew-product case [17, Corollary 1.4]. Another natural future direction is to derive a criterion for uniqueness. In [17] (see also [18]) it is shown that each homoclinic class supports a unique measure of maximal entropy. In particular, any smooth, transitive surface system with positive entropy has a unique measure of maximal entropy. The analogous statement for measures of relative maximal entropy remains unclear.
It is worth mentioning that, as a consequence of a celebrated result due to Newhouse [34], $\mathcal{C}^{\infty}$ skew products always admit a measure of maximal $\mathbb{P}$-relative entropy.
Proposition 1.5.
Assume that the skew product $\Theta$ is $\mathcal{C}^{\infty},$ and let $\mathbb{P}$ be any $\theta$-invariant measure. Then $\Theta$ admits a $\mathbb{P}$-relative measure of maximal entropy.
Proof.
From [34, Theorem 4.1] the map
$\mu\in\left\{\nu;\ \nu\text{ is a $\Theta$-invariant probability measure}\right\}\mapsto h_{\mu}(\Theta)\in\mathbb{R}$
is upper semi-continuous with respect to the weak topology. Since
$\mathcal{M}_{\mathbb{P}}(\Theta):=\left\{\nu;\ \nu\text{ is }\Theta\text{-invariant probability measure }\text{ and }\nu(\mathrm{proj}_{\Omega}^{-1}(\cdot))=\mathbb{P}\right\}$
is compact in the weak topology, there exists $\widetilde{\mu}\in\mathcal{M}_{\mathbb{P}}(\Theta)$ such that
$h_{\widetilde{\mu}}(\Theta)=\sup\left\{h_{\nu}(\Theta);\ \nu\in\mathcal{M}_{\mathbb{P}}(\Theta)\right\}.$
Since $h_{\nu}(\Theta\mid\mathbb{P})=h_{\nu}(\Theta)-h_{\mathbb{P}}(\theta)$ for each $\nu\in\mathcal{M}_{\mathbb{P}}(\Theta)$ we obtain that $\widetilde{\mu}$ is a measure of maximal $\mathbb{P}$-relative entropy. ∎
Throughout the paper, if $X$ is a topological space and $T:X\to X$ is a measurable map, we denote:
  • $\mathcal{B}\left(X\right)$ as the Borel $\sigma$-algebra of $X$;
  • $\mathcal{M}(X):=\left\{\mu;\ \mu\ \text{is a positive measure on }X\right\};$
  • $\mathcal{M}(T):=\left\{\mu\in\mathcal{M}\left(X\right);\ T_{*}\mu:=\mu(T^{-1}(\cdot))=\mu\ \text{and }\mu\ \text{is a probability measure}\right\};$
  • $\mathcal{M}_{e}(T):=\{\mu\in\mathcal{M}(T);\ \mu\ \text{is ergodic}\}$.
1.3. An application of Corollary B
Let $\mathbb{T}^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$ be the flat torus. A classical and notoriously challenging dynamical system is the standard map on $\mathbb{T}^{2}$. One convenient version of the standard map is $T:\mathbb{T}^{2}\to\mathbb{T}^{2}$ given by
$T(x_{1},x_{2})=(2x_{1}-x_{2}+k\cos(2\pi x_{1}),\,x_{1})\pmod{1},\ \text{for some }k\in\mathbb{Z}.$
Sinai conjectured that for $k$ sufficiently large, $T$ has positive metric entropy with respect to Lebesgue measure [42, Page 144]. This conjecture remains open. In contrast, much more is known about measures of maximal entropy: Obata showed that, for large $k$, the standard map admits a unique measure of maximal entropy; moreover, such a measure has positive metric entropy [35]. However, less is known in the random setting. Our results show that there exists at least one and, at most, countably many ergodic relative measures of maximal entropy for certain classes of random perturbations of the standard map.
For each $k\in\mathbb{N}$, we consider the skew product
$\Theta_{k}:\mathbb{T}^{2}\times\mathbb{T}^{2}$ $\to\mathbb{T}^{2}\times\mathbb{T}^{2}$
$(\omega,x)$ $\mapsto(A\omega,T_{\omega}(x)),$
Where $A:\mathbb{T}^{2}\to\mathbb{T}^{2}$ is Arnold’s cat map
$A(\omega_{1},\omega_{2})=(2\omega_{1}+\omega_{2},\omega_{1}+\omega_{2})\pmod{1}$,
and $T_{\omega}:\mathbb{T}^{2}\to\mathbb{T}^{2}$ is a perturbed version of the standard map
$T_{\omega}(x_{1},x_{2}):=(2x_{1}-x_{2}+k\cos(2\pi x_{1})+f(\omega),x_{1})\pmod{1},\ \text{for every }\omega\in\Omega,$
where $f:\mathbb{T}^{2}\to\mathbb{S}^{1}$ is a $\mathcal{C}^{\infty}$ function.
Proposition 1.6.
Let $\mathbb{P}$ be an ergodic $\theta$-invariant measure. Then, there exists $k_{0}\in\mathbb{N}$, such that for every $k\geq k_{0},$ one has $h_{\mathrm{top}}(\Theta_{k}\mid\mathbb{P})>0$. In particular, Proposition 1.5 implies that $\Theta_{k}$ admits a measure of maximal $\mathbb{P}$-relative entropy and Corollary B ensures that there are only countably many such ergodic measures.
Remark 1.7.
Examples of choices for $f$ are:
  • $f(\omega_{1},\omega_{2})=\omega_{1}$ (large random additive kicks);
  • $f(\omega_{1},\omega_{2})=\varepsilon\sin(2\pi\omega_{1})$ for any $\varepsilon>0$ (arbitrarily small random additive kicks); and
  • $f(\omega_{1},\omega_{2})=2\omega_{1}+\varepsilon\sin(2\pi\omega_{2})$ (random kicks of mixed sizes depending on both base coordinates).
2. Hyperbolic measures and Borel homoclinic classes
In this section, we collect standard facts from hyperbolic dynamics and introduce the concept of Borel homoclinic classes (see [18, Definition 2.7]). Fix a compact smooth manifold $E$ and a $\mathcal{C}^{1+\alpha}$ diffeomorphism $\Theta:E\to E$, for some $\alpha>0$. Unless stated otherwise, all statements in this section are made under this standing assumption. For the purposes of this paper, the set $E$ can be thought of as $\Omega\times M$ as in Theorem A. Our exposition closely follows [18, Section 2].
2.1. Hyperbolic measures
By the Oseledets’ multiplicative ergodic theorem [36], there exists a Borel set $E^{\prime}\subset E$ such that
  • $E^{\prime}$ has full measure for every $\Theta$-invariant probability measure; and
  • for every $x\in E^{\prime}$ there is a finite set $\sigma(x)\subset\mathbb{R}$ and a $\mathrm{D}\Theta$-invariant splitting
    $T_{x}E=\bigoplus_{\lambda\in\sigma(x)}E_{x}^{\lambda},$
    where each $E_{x}^{\lambda}\neq\{0\}$ is a linear subspace defined by
    $E_{x}^{\lambda}=\left\{v\in T_{x}E\setminus\{0\};\ \lim_{n\to\infty}\frac{1}{n}\log\big\|\mathrm{D}_{x}\Theta^{n}v\big\|=\lambda\right\}\ \cup\ \{0\}.$
Definition 2.1.
Let $\Theta:E\to E$ be as above and $x\in E^{\prime}$. The set
$\sigma(x)=\{\lambda_{1}(x),\ldots,\lambda_{k(x)}(x)\}$
is called the Lyapunov spectrum of $\Theta$ at $x$, its elements are the Lyapunov exponents of $\Theta$ at $x$. For $\chi>0$ and $\mu\in\mathcal{M}(\Theta)$ we say that $\mu$ is $\chi$-hyperbolic if, for $\mu$-a.e. $x\in E$.
  • (i)
    $\min\{|\lambda|;\ \lambda\in\sigma(x)\}\geq\chi$;
  • (ii)
    there are both positive and negative exponents, i.e.
    $\sigma^{+}(x):=\{\lambda\in\sigma(x):\lambda>0\}\neq\ \emptyset\quad\text{and}\ \sigma^{-}(x):=\{\lambda\in\sigma(x):\lambda<0\}\neq\emptyset.$
We then set
$E_{x}^{s}:=\bigoplus_{\lambda\in\sigma^{-}(x)}E_{x}^{\lambda}\ \text{and}\ E_{x}^{u}:=\bigoplus_{\lambda\in\sigma^{+}(x)}E_{x}^{\lambda}.$
If $\mu\in\mathcal{M}_{e}(\Theta)$ is ergodic, then there exist real numbers $\lambda_{1}<\cdots<\lambda_{k}$ and integers $m_{1},\ldots,m_{k}\geq 1$ such that, for $\mu$-a.e. $x\in E$,
$\sigma(x)=\{\lambda_{1},\ldots,\lambda_{k}\}\ \text{and}\ \dim E_{x}^{\lambda_{i}}=m_{i},\text{ for }i=1,\ldots,k.$
Finally, we recall that if (i) holds but (ii) fails, then $\mu$ is supported on a periodic orbit and hence has zero metric entropy [26, Corollary S.5.2] (see also the proof of [25, Theorem 4.2]).
Below, we recall a few results of Pesin theory. These definitions will be useful for defining the concept of Borel homoclinic classes in the following section.
Definition 2.2 ($(\chi,\varepsilon)$-Pesin block).
Given $\chi,\varepsilon>0$, a $(\chi,\varepsilon)$-Pesin block is a non-empty set $\Lambda\subset E$ for which there exists a direct sum $T_{x}E=E^{s}(x)\oplus E^{u}(x)$ for all $x\in\bigcup_{n\in\mathbb{Z}}\Theta^{n}(\Lambda),$ and a number $K>0$ such that for any $n\in\mathbb{Z}$, $k\geq 0,$ and $y\in\Lambda$,
$\mathrm{max}\left(\left\|\left.\mathrm{D}\Theta^{k}\right|_{E^{s}(\Theta^{n}(y))}\right\|,\left\|\left.\mathrm{D}\Theta^{-k}\right|_{E^{u}(\Theta^{n}(y))}\right\|\right)\leq Ke^{-\chi k+\varepsilon|n|}.$
We recall that Pesin blocks are not necessarily $\Theta$-invariant sets.
Below we state Pesin’s stable manifold theorem, whose proof can be found at [38] (see also [18, Theorem 2.2]).
Theorem 2.3.
There exists a continuous function $\varepsilon_{0}$ with the following property: Given $\chi,\varepsilon>0$, with $\varepsilon<\varepsilon_{0}(\chi)$, and a $(\chi,\varepsilon)$-Pesin block $\Lambda$. Then each $x\in\Lambda$ admits a $\mathcal{C}^{1+\alpha}$ embedded disc $W_{\mathrm{loc}}^{s}(x)$ satisfying:
$\text{for each }y\in W^{s}_{\mathrm{loc}}(x),\ \limsup_{n\to\infty}\frac{1}{n}\log\mathrm{dist}_{N}(\Theta^{n}(x),\Theta^{n}(y))<0.$
Moreover, the embedded disc $W_{\mathrm{loc}}^{s}$ varies continuously with $x\in\Lambda$ in the $\mathcal{C}^{1+\alpha}$ topology. The same result applied to $\Theta^{-1}$ gives the local unstable manifold, which we call $W^{u}_{\mathrm{loc}}(x).$
The above theorem allows us to define the global stable and unstable manifolds of $x$ lying in a $(\chi,\varepsilon)$-Pesin block, where $(\chi,\varepsilon)$ satisfy the assumptions of Theorem 2.3, as
$W^{s}(x):=\bigcup_{n\geq 0}\Theta^{-n}(W^{s}_{\mathrm{loc}}(\Theta^{n}(x)))\ \text{and }W^{u}(x)=\bigcup_{n\geq 0}\Theta^{n}(W^{s}_{\mathrm{loc}}(\Theta^{-n}(x))).$
We recall that the choice of $\varepsilon>0$ in Theorem 2.3 does not affect the definition of the global stable and unstable manifolds [8, §8.2] (see also [18, Section 2]).
Definition 2.4.
We define the recurrent non-uniformly hyperbolic set of $\Theta$ as
[…]
Observe that
$\mathrm{NUH^{\prime}}(\Theta)\subset\bigcup_{\chi>0}\bigcap_{\varepsilon>0}\{x\in E:\,x\ \text{is contained in a }(\chi,\varepsilon)\text{-Pesin block}\}.$
In particular, from Theorem 2.3 each $x\in\mathrm{NUH}^{\prime}(\Theta)$ admits global stable and unstable manifolds.
Theorem 2.5 ([25, Theorem 4.1]).
Let $\mu\in\mathcal{M}_{e}(\Theta)$, then $\mu(\mathrm{NUH}^{\prime}(\Theta))=1$.
2.2. Borel homoclinic classes
Given two submanifolds $S_{1},S_{2}$ in $E$ we denote $S_{1}\pitchfork S_{2}$ to be the set of transverse intersection points, i.e.
$S_{1}\pitchfork S_{2}:=\{x\in S_{1}\cap S_{2}:T_{x}S_{1}\oplus T_{x}S_{2}=T_{x}E\}.$
Definition 2.6 ([18], Homoclinic relation).
Two points $x,y\in\mathrm{NUH}^{\prime}(\Theta)$ are homoclinically related $(x\sim y)$ if $W^{s}(x)$ has a transverse intersection point with an iterate of $W^{u}(y)$, and $W^{u}(x)$ has a transverse intersection with an iterate of $W^{s}(y)$.
Proposition 2.7 ([18, Proposition 2.6]).
The homoclinic relation $\sim$ is an equivalence relation on $\mathrm{NUH}^{\prime}(\Theta)$.
Definition 2.8 ([18]).
Borel homoclinic classes are equivalence classes of $\sim$ in $\mathrm{NUH}^{\prime}(\Theta)$.
The proposition below states that the homoclinic classes characterise all the hyperbolic measures.
Proposition 2.9 ([18, Proposition 2.8]).
The following hold:
  1. (i)
    The set of Borel homoclinic classes is a finite or countable partition of $\mathrm{NUH}^{\prime}(\Theta)$ into invariant Borel sets;
  2. (ii)
    each class contains a dense set of hyperbolic periodic points;
  3. (iii)
    any hyperbolic ergodic measure $\mu$ is carried by a Borel homoclinic class $\mathcal{X}$; and
  4. (iv)
    any invariant measure carried by a Borel homoclinic class is hyperbolic.
Remark 2.10.
Note that any Borel homoclinic class that is not a singleton supports uncountably many ergodic invariant measures. If $\Theta$ admits a non-atomic hyperbolic invariant measure (i.e., not supported on a periodic orbit), then $\Theta$ admits uncountably many hyperbolic ergodic measures. Indeed, every non-trivial Borel homoclinic class contains at least one horseshoe, and each horseshoe carries uncountably many ergodic measures since it is conjugated to a full shift. By Proposition 2.9 (iv), each of these measures is hyperbolic.
3. Topological Markov shifts
This section proves Theorem 3.5, which enables us to analyse the skew product $\Theta$ via a countable-state topological Markov shift and a factor map that preserves the base dynamics.
Definition 3.1.
Let $S$ be a countable alphabet and a directed graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$, such that the set of vertices $\mathscr{V}=S$. The topological Markov shift associated to $\mathscr{G}$ is the discrete-time topological dynamical system $\sigma:\Sigma\to\Sigma$ where
$\Sigma=\Sigma(\mathscr{G})=\{v=(v_{i})_{i\in\mathbb{Z}}\in S^{\mathbb{Z}};\ \text{if }v_{i}\to v_{i+1}\in\mathscr{E}\ \text{for every }i\in\mathbb{Z}\},$
equipped with the metric $d(v,u)=\min\{\exp(-\min\{|n|;v_{n}\neq u_{n}\})\}$, and $\sigma:\Sigma\to\Sigma$ is the left shift map $\sigma((v_{i})_{i\in\mathbb{Z}})=(v_{i+1})_{i+1}$. Given a topological Markov shift $\Sigma$ associated to a graph $\mathrm{G}$ we denote as $V(\Sigma)$ the alphabet of $\Sigma$ and we define the set
$\Sigma^{\#}=\left\{(v_{i})_{i\in\mathbb{Z}}\in\Sigma;\,\exists u^{\prime},v^{\prime}\in V(G)\ \text{s.t. }\,\#\{i\in\mathbb{Z}_{+};v_{i}=v^{\prime}\}=\#\{i\in\mathbb{Z}_{-};v_{i}=u^{\prime}\}=\infty\right\}.$ (2)
We say that $\Sigma$ is irreducible if for every two symbols $v_{i},v_{j}\in\mathscr{V}$, there exists a path in the graph $\mathscr{G}$ which connects $v_{i}$ to $v_{j}$.
Definition 3.2 (Bowen Partition).
Let $\theta:\Omega\to\Omega$ be a homeomorphism. We say that $\theta$ admits a Bowen partition if there exist a family of sets $\mathscr{P}\subset\mathcal{B}(\Omega)$ such that
  • (1)
    $\mathscr{P}<\infty$;
  • (2)
    let $\mathscr{G}_{b}=(\mathscr{V}_{b},\mathscr{E}_{b})$, with $\mathscr{V}_{b}=\mathscr{P}$ and
    $\mathscr{E}_{b}=\{P\to P^{\prime};\,P,P^{\prime}\in\mathscr{P}\ \text{and }P\cap\theta^{-1}(P^{\prime})\neq\emptyset\}.$
    Setting the subshift of finite type $\Sigma_{b}=\Sigma_{b}(\mathscr{G}_{b})$ the have that map
    $\pi_{b}:(P_{i})_{i\in\mathbb{Z}}\in\Sigma_{b}\mapsto\bigcap_{i\in\mathbb{Z}}\theta^{-i}(P_{i})\in\Omega,$ (3)
    is a well-defined, Hölder continuous and each $\pi^{-1}(\omega)\leq k_{\mathscr{\mathscr{P}}}$ for some $k_{\mathscr{P}}\in\mathbb{P}$; and
  • (3)
    the function $\pi$ is surjective over the non-wandering set $NW(\theta)$ and there exists $k<\infty$ such that $\pi^{-1}(\omega)\leq k$ for every $\omega\in NW(\theta)$.
Observe that under the above assumptions imply that $\pi\circ\sigma=\theta\circ\pi.$
Theorem 3.3 ([13], [37, Appendix III]).
Every $\mathcal{C}^{1+\alpha}$Axiom $A$ diffeomorphism admits a Bowen partition.
The next statement, taken from [18, Theorem 10.5 and Section 10], is a higher-dimensional generalisation of Sarig’s celebrated theorem [40, Theorem 1.3] (see also [9, Theorem 0.1]). We mention that the original result presented in [18] is more general; however, the presented version is enough for our purposes.
Theorem 3.4 ([18, Theorem 10.5 and Section 10]).
Let $\Theta:E\to E$ be a $\mathcal{C}^{1+\alpha}$ diffeomorphism on a compact smooth manifold $E$. Let $\mathcal{X}$ be either a Borel homoclinic class $H$ or $E$. Then, for each $\chi>0$, there exists a countable collection of sets $\mathscr{R}\subset\mathscr{B}(M)$, such that the topological Markov shift $\Sigma=\Sigma(\mathscr{G})$, induced by the graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$ with $\mathscr{V}:=\mathscr{R}$ and $\mathscr{E}=\{R\to R^{\prime};R,R^{\prime}\in\mathscr{R}\ \text{and }R\cap\Theta^{-1}(R^{\prime})\neq\emptyset\},$ satisfies the following properties:
  1. (1)
    The map
    $\pi:(R_{i})_{i\in\mathbb{Z}}\in\Sigma\mapsto\bigcap_{i\in\mathbb{Z}}\Theta^{-1}(R_{i})\in\mathcal{X}$
    is well defined and Hölder continuous. Moreover, $f\circ\pi=\pi\circ\sigma$ and $\pi:\Sigma^{\#}\to\mathcal{X}$ is finite-to-one (see Definition 3.1).
  2. (2)
    $\Sigma$ is a locally compact countable-state Markov shift and $\mu(\pi(\Sigma))=\mu(\pi(\Sigma^{\#}))=1$. If $\mathcal{X}$ is a Borel homoclinic class, we can choose $\Sigma$ as irreducible;
  3. (3)
    Every ergodic invariant probability measure $\widehat{\mu}$ on $\Sigma$ projects to an ergodic $\Theta$-invariant measure $\mu:=\pi_{*}\widehat{\mu}$ such that $\mu(\mathcal{X})=1$ and $h_{\mu}(\Theta)=h_{\widehat{\mu}}(\sigma).$
  4. (4)
    For every $\chi$-hyperbolic $\Theta$-invariant probability measure $\mu$ on $M$ there exists a $\sigma$-invariant probability measure $\widehat{\mu}$ on $\Sigma$ such that $\pi_{*}\widehat{\mu}=\mu$.
We call $\mathscr{R}$ a $\chi$-Sarig partition of $\Theta$.
We will prove the following theorem:
Our results are based on the following theorem, which is based on a refinement of Theorem 3.4 for skew products.
Theorem 3.5.
Let $\Theta:\Omega\times M\to\Omega\times M$ be a $\mathcal{C}^{1+\alpha}$ diffeomorphism such that
  1. (a)
    $\Theta$ is a skew product i.e. $\Theta(\omega,x)=(\theta(\omega),T_{\omega}(x))$; and
  2. (b)
    $\theta:\Omega\to\Omega$ is Axiom $A$.
Let $\mathcal{X}\subset\Omega\times M$ be equal to a Borel homoclinic class $H$ or $\Omega\times M$. Then, for every $\chi>0$ there exist a locally compact topological Markov shift $\Sigma_{s}$, a subshift of finite type $\Sigma_{b}$ and Hölder continuous maps $\pi_{s}:\Sigma_{s}\to\mathcal{X}$ and $\pi_{b}:\Sigma_{b}\to\Omega$ such that
  1. (1)
    $\pi_{s}\circ\sigma_{s}=\Theta\circ\pi_{s}$ and $\pi_{b}\circ\sigma_{b}=\theta\circ\pi_{b}$, where $\sigma_{s}$ is the left shift map on $\Sigma_{s}$ and $\sigma_{b}$ is the left shift map on $\Sigma_{b}$;
  2. (2)
    for every $\chi$-hyperbolic $\Theta$-invariant probability $\mu$ such that $\mu(\mathcal{X})=1$, one has $\mu\big(\pi_{s}[\Sigma_{s}]\big)=\mu\big(\pi_{s}[\Sigma_{s}^{\#}]\big)=1$};
  3. (3)
    the maps $\pi_{s}|_{\Sigma_{s}^{\#}}:\Sigma_{s}^{\#}\to\mathcal{X}$ and $\pi_{b}:\Sigma_{b}\to\Omega$ are finite-to-one;
  4. (4)
    there exists a one $1$-block factor map $\mathrm{proj}_{\Sigma_{b}}:\Sigma_{s}\to\Sigma_{b}$ such that
    $\pi_{b}\circ\mathrm{proj}_{\Sigma_{b}}=\mathrm{proj}_{\Omega}\circ\pi_{s}.$
    Write $\Sigma_{b}^{\prime}:=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})$. If $\mathcal{X}=\Omega\times M$ then $\mathrm{proj}_{\Sigma_{b}}$ is onto and $\Sigma_{b}^{\prime}=\Sigma_{b}$. If $\mathcal{X}$ is a Borel homoclinic class, then $\Sigma_{b}^{\prime}$ is a subshift of finite type; and
  5. (5)
    if $\mathcal{X}$ is a Borel homoclinic class, $\Sigma_{s}$ can be chosen irreducible; in this case $\Sigma_{b}^{\prime}$ is an irreducible subshift of finite type.
Diagram 1 summarises all of the sets and projections used in the following proof.
$\mathcal{X}\subset\Omega\times M$$\mathcal{X}\subset\Omega\times M$$\Sigma$$\Sigma$$\Sigma_{s}$$\Sigma_{s}$$\Sigma_{b}^{\prime}\subset\Sigma_{b}$$\Sigma_{b}^{\prime}\subset\Sigma_{b}$$NW(\theta)\subset\Omega$$NW(\theta)\subset\Omega$$\scriptstyle{{\rm proj}_{\Omega}}$$\scriptstyle{\Theta}$$\scriptstyle{{\rm proj}_{\Omega}}$$\scriptstyle{\sigma}$$\scriptstyle{\pi}$$\scriptstyle{\pi}$$\scriptstyle{\pi_{s}}$$\scriptstyle{{\rm proj}_{\Sigma_{b}}}$$\scriptstyle{\sigma_{s}=\sigma_{b}\times\sigma}$$\scriptstyle{{\rm proj}_{\Sigma}}$$\scriptstyle{{\rm proj}_{\Sigma}}$$\scriptstyle{\pi_{s}}$$\scriptstyle{{\rm proj}_{\Sigma_{b}}}$$\scriptstyle{\sigma_{b}}$$\scriptstyle{\pi_{b}}$$\scriptstyle{\pi_{b}}$$\scriptstyle{\theta}$
Diagram 1: Commutative diagram of the maps $\Theta,\pi,\pi_{s},\pi_{b},\mathrm{proj}_{\Sigma},\mathrm{proj}_{\Sigma_{b}}$, and $\mathrm{proj}_{\Omega}$. We note that $\Sigma_{s}\subsetneq\Sigma_{b}\times\Sigma,$ in fact, $\Sigma_{s}=\{\underline{(P,R)};\underline{P}\in\Sigma_{s},\,\underline{R}\in\Sigma\ \text{and }\pi_{b}(\underline{P})=\mathrm{proj}_{\Omega}\circ\pi(\underline{R})\}$.
Proof of Theorem 3.5.
Let $\chi>0$. Applying Theorem 3.4 to the skew product $\Theta:\Omega\times M\to\Omega\times M$ and $\mathcal{X}$ be either equal to Borel homoclinic class $H$ or $\Omega\times M$, we obtain a topological Markov shift $(\Sigma,\sigma)$ with the properties asserted therein, and let $\mathscr{R}$ be its $\chi$-Sarig partition. We assume $(\Sigma,\sigma)$ to be irreducible if $\mathcal{X}$ is a Borel homoclinic class. Since $\theta:\Omega\to\Omega$ is an Axiom A diffeomorphism, there exists a subshift of finite type $(\Sigma_{b},\sigma_{b})$ associated with a Bowen partition $\mathscr{P}$ as in Definition 3.2, let $\pi_{b}$ be as in (3) and recall that
$k_{\mathscr{P}}:=\sup_{\omega\in NW(\theta)}\#\pi_{b}^{-1}(\omega)<\infty.$
Define the graph $\mathscr{G}_{s}:=(\mathscr{V}_{s},\mathscr{E}_{s})$ where
$\mathscr{V}_{s}:=\{(P,R);\,R\in\mathscr{R},P\in\mathscr{P}\ \text{and }R\cap(P\times M)\neq\emptyset\},$
and
$\mathscr{E}_{s}:=\left\{(P,R)\to(P^{\prime},R^{\prime});\,\Theta^{-1}(R^{\prime})\cap R\neq\emptyset\ \text{and }P^{\prime}\cap\theta^{-1}(P)\neq\emptyset\right\}.$
Define $(\Sigma_{s},\sigma_{s}):=(\Sigma_{s}(\mathscr{G}_{s}),\sigma_{s})$ and consider the map
$\pi_{s}:((P_{i},R_{i}))_{i\in\mathbb{Z}}\in\Sigma_{s}\mapsto\bigcap_{i\in\mathbb{Z}}R_{i}\in\Omega\times M.$
We divide the remaining of the proof into five steps.
Step 1.
We show item $(1).$
Observe that the map $\pi_{s}$, defined above, is well defined and Hölder continuous. Indeed
$\mathrm{dist}_{\Omega\times M}(\pi_{s}\left(\{(P_{i},R_{i})\}_{i\in\mathbb{Z}}\right),\pi_{s}\left(\{(P_{i}^{\prime},R_{i}^{\prime})\}_{i\in\mathbb{Z}}\right)$ $\leq\mathrm{dist}_{\Omega\times M}\left(\pi\left((R_{i})_{i\in\mathbb{Z}}\right),\pi\left((R_{i}^{\prime})_{i\in\mathbb{Z}}\right)\right)$
$\leq C\mathrm{dist}_{\Sigma}((R_{i})_{i\in\mathbb{Z}},(R_{i}^{\prime})_{i\in\mathbb{Z}})^{\eta}$
$\leq C\mathrm{dist}_{\Sigma_{s}}(\{(P_{i},R_{i})\}_{i\in\mathbb{Z}},\{(P_{i}^{\prime},R_{i}^{\prime})\}_{i\in\mathbb{Z}})^{\eta},$
where $C$ and $\eta$ are the Hölder map constants of $\pi:\Sigma\to\Omega\times M$. Moreover, it is clear from the definitions of $\pi_{s}$ and $\pi_{b}$ that
$\Theta\circ\pi_{s}=\pi_{s}\circ\sigma_{s}\ \text{and }\theta\circ\pi_{b}=\pi_{b}\circ\sigma_{b}.$ (4)
Item $(1)$ now follows.
Step 2.
We show that if $\mu$ is a $\chi$-hyperbolic $\Theta$-invariant measure such that $\mu(\mathcal{X})=1$ then $\mu(\pi_{s}(\Sigma_{s}))=\mu(\pi_{s}(\Sigma_{s}^{\#}))=1,$ which implies item (2).
Recall from Theorem 3.4 (2) that if $\mu(\mathcal{X})=1$, then $\mu(\pi(\Sigma))=\mu(\pi(\Sigma^{\#}))=1$ for each $\chi$-hyperbolic $\Theta$-invariant measure. Since $\Sigma_{s}^{\#}\subset\Sigma_{s}$, it is enough to show that for each $\underline{R}=(R_{i})_{i\in\mathbb{Z}}\in\Sigma^{\#}$ there exists $\underline{P}=(P_{i})_{i\in\mathbb{Z}}\in\Sigma_{b}$ such that $\underline{(P,R)}=\{(P_{i},R_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{s}^{\#}.$ Indeed, under this assumption we obtain
$\mu(\pi_{s}(\Sigma))\geq\mu(\pi_{s}(\Sigma^{\#}))\geq\mu(\pi(\Sigma^{\#}))=1$
as required.
First of all, observe that if $(\omega,x)=\pi(\underline{R})$. For each $i\in\mathbb{Z}$ let $P_{i}\in\mathscr{P}$ such that $\theta^{i}(\omega)\in P_{i}$. It is clear that $\underline{P}=(P_{i})_{i\in\mathbb{Z}}\in\Sigma_{b}$. Moreover, $\Theta^{i}(\omega,x)=(\theta\omega,T_{\omega}(x))\in R_{i}$ for each $i\in\mathbb{Z}$. In this way,
$R_{i}\cap(P_{i}\times M)\neq\emptyset,\,\Theta^{-1}(R_{i+1})\cap R_{i}\neq\emptyset\ \text{and}\ \theta^{-1}(P_{i+1})\cap P_{i}\neq\emptyset\ \text{for every }i\in\mathbb{Z}.$
Therefore $\underline{(P,R)}\in\Sigma_{s}$ and $\pi_{s}(\underline{(P,R)})=\pi(\underline{R})$, in the following we ensure that $\underline{(P,R)}\in\Sigma_{s}^{\#}$.
Recall that $\underline{R}\in\Sigma^{\#}$. Since $\underline{R}\in\Sigma^{\#}$ there exist $R^{+},R^{-}\in\mathscr{R}$ and increasing sequences of natural numbers $\{i_{k}\}_{k\in\mathbb{N}},\{j_{k}\}_{\in\mathbb{N}}\subset\mathbb{N}$ such that
$R_{i_{k}}=R^{+}\ \text{and }R_{-j_{k}}=R^{-}\ \text{for each }k\in\mathbb{N}.$
Recall that $\underline{P}\in\Sigma_{b}$ and $\Sigma_{b}$ is a shift which admits finitely many symbols. Then, there exists $P^{+},P^{-}\in\mathscr{P}$ and subsequences $\{i_{k}^{\prime}\}_{k\in\mathbb{N}}\subset\{i_{k}\}_{k\in\mathbb{N}}$ and $\{j_{k}^{\prime}\}_{\in\mathbb{N}}\subset\{j_{k}\}_{\in\mathbb{N}}$ such that
$(P,R)_{i_{k}^{\prime}}=(P_{i_{k}^{\prime}},R_{i_{k}^{\prime}})=(P^{+},R^{+})\ \text{and }(P,R)_{j_{k}^{\prime}}=(P_{-j_{k}^{\prime}},R_{-j_{k}^{\prime}})=(P^{-},R^{-})\ \text{for each }k\in\mathbb{N}.$
Thus, $\underline{(P,R)}\in\Sigma^{\#}_{s}$ and $\pi_{s}(\underline{(P,R)})=\pi(\underline{R}).$ Which concludes the proof of Step 2.
Step 3.
We show that the Lipschitz 1-block map
$\mathrm{proj}_{\Sigma}:\ \Sigma_{s}\to\Sigma,\quad\big((P_{i},R_{i})_{i\in\mathbb{Z}}\big)\mapsto(R_{i})_{i\in\mathbb{Z}},$
satisfies $\mathrm{proj}_{\Sigma}^{-1}(\underline{R})\leq k_{\mathscr{P}}$ for every $\underline{R}\in\Sigma$, and item $(3)$.
First of all, observe that given $\underline{(P,R)},\underline{(P^{\prime},R^{\prime})}\in\Sigma_{s}$ we have that
$\mathrm{dist}_{\Sigma}(\mathrm{proj}_{\Sigma}(\underline{(P,R}),\mathrm{proj}_{\Sigma}(\underline{(P^{\prime},R^{\prime}})))=\mathrm{dist}_{\Sigma}(\underline{R},\underline{R^{\prime}})\leq\mathrm{dist}_{\Sigma}(\underline{(P,R)},\underline{(P^{\prime},R^{\prime})})$ (5)
Let $(\omega,x)=\pi(\underline{R})$. Recall that $\pi_{b}:\Sigma_{b}\to\Omega$ is a finite-to-one surjective map over $NW(\theta)$. Therefore $\pi_{b}^{-1}(\omega)=\{\underline{P}^{(1)},\ldots,\underline{P}^{(k)}\}$ for some $k\in\mathbb{N}$. To conclude the first part of Step 3, observe that:
$\#\{\mathrm{proj}_{\Sigma}^{-1}(\underline{R})\}$ $=\#\left\{\underline{(P,R)}\in\Sigma_{s};\underline{P}\in\Sigma_{b}\right\}$
$=\#\left\{\underline{(P,R)}\in\Sigma_{s};\theta^{i}\omega\in P_{i}\ \text{and }R_{i}\cap(P_{i}\times M)\neq\emptyset\ \text{for each }i\in\mathbb{Z}\right\}$
$\leq\#\left\{\underline{P}\in\Sigma_{b};\pi_{b}(\underline{P})=\omega\right\}=\#\pi_{b}^{-1}(\omega)\leq k_{\mathscr{P}}.$
We claim that $\underline{(P,R)}\in\Sigma_{s}^{\#}$ if, and only if, $\underline{R}=\mathrm{proj}_{\Sigma}(\underline{(P,R)})\in\Sigma^{\#}.$ Indeed, if $\underline{(P,R)}\in\Sigma_{s}^{\#}$, then by (2) there exist $(P_{1},R_{1}),(P_{2},R_{2})\in\mathscr{V}_{s}$, such that
$\{i\in\mathbb{N};(P,R)_{i}=(P_{1},R_{1})\}=\#\{i\in\mathbb{N};(P,R)_{-i}=(P_{2},R_{2})\}=\infty.$
Therefore $\underline{R}\in\Sigma^{\#}.$ The proof of the reverse implication can be found in the last paragraph of Step 2. Hence,
$\left\{\left(\pi_{s}|_{\Sigma_{s}^{\#}}\right)^{-1}(\omega,x)\right\}=\#\left\{\mathrm{proj}_{\Sigma}^{-1}\left(\left(\pi|_{\Sigma^{\#}}\right)^{-1}(\omega,x)\right)\right\}.$
Since the maps $\mathrm{proj}_{\Sigma}$ and $\left.\pi\right|_{\Sigma^{\#}}$ are both finite-to-one, we obtain that
$\left\{\left(\pi_{s}|_{\Sigma_{s}^{\#}}\right)^{-1}(\omega,x)\right\}<\infty.$
Hence $\pi_{s}:\Sigma_{s}^{\#}\to\mathcal{X}$ has uniformly bounded fibres and is therefore finite-to-one. Independently, since $\pi_{b}:\Sigma_{b}\to\Omega$ is obtained from a Bowen partition for $\theta$, it is finite-to-one on $NW(\theta)$. This proves item (3).
Step 4.
We show item (4).
Define
$\mathrm{proj}_{\Sigma_{b}}:\{(P_{i},R_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{s}\mapsto\{(P_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{b},$
and set $\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s}).$ Observe that $\mathrm{proj}_{\Sigma_{b}}$ is a Lipschitz $1$-block factor map, in fact:
$\mathrm{dist}_{\Sigma_{b}}(\mathrm{proj}_{\Sigma_{b}}(\underline{(P,R}),\mathrm{proj}_{\Sigma_{b}}(\underline{(P^{\prime},R^{\prime}})))=\mathrm{dist}_{\Sigma_{b}}(\underline{P},\underline{P^{\prime}})\leq\mathrm{dist}_{\Sigma}(\underline{(P,R)},\underline{(P^{\prime},R^{\prime})}).$ (6)
Let $\underline{(P,R)}\in\Sigma_{s}$ and $\pi_{s}(\underline{(P,R)})=(\omega,x)$. Therefore,
$\Theta^{i}(\omega,x)=(\theta^{i}(\omega),T_{\omega}^{i}(x))\in R_{i}\ \text{and }R_{i}\cap(P_{i}\times M)\neq\emptyset\text{ for every }i\in\mathbb{N}.$
In particular $\theta^{i}(\omega)\in P_{i}$ for each $i\in\mathbb{Z}$ and therefore $\pi_{b}(\underline{P})=\omega.$ Hence
$\mathrm{proj}_{\Omega}\circ\pi_{s}(\underline{(P,R)})=\mathrm{proj}_{\Omega}(\omega,x)=\omega=\pi_{b}(\underline{P})=\pi_{b}\circ\mathrm{proj}_{\Sigma_{b}}(\underline{(P,R)}).$
Observe that $\mathrm{proj}_{\Sigma_{b}}$ is clearly surjective if $\mathcal{X}=\Omega\times M$, therefore $\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})=\Sigma_{b}$. If $\mathcal{X}$ is a Borel homoclinic class, observe that, by construction
$\Sigma_{b}^{\prime}=\Sigma(\mathscr{G}^{\prime}),$ (7)
where $\mathscr{G}^{\prime}=(\mathscr{V}^{\prime},\mathscr{E}^{\prime})$ with
$\mathscr{V}^{\prime}=\{P\in\mathcal{P}^{\prime};\text{there exists }R\in\mathscr{R}\ \text{such that }(P\times M)\cap R\neq\emptyset\},$
and
$\mathscr{E}^{\prime}=\left\{P\to P^{\prime}\;\middle|\;\begin{aligned} &P,P^{\prime}\in\mathscr{V}^{\prime},\ \theta^{-1}(P^{\prime})\cap P\neq\emptyset\ \text{and there exist }R,R^{\prime}\in\mathscr{R},\\ &\ \text{such that }\Theta^{-1}(R^{\prime}\cap(P^{\prime}\times M))\cap(R\cap(P\times M))\neq\emptyset\end{aligned}\right\}.$
In this way, $\Sigma_{b}^{\prime}$ is a subshift of finite type.
Step 5.
We show item (5) and therefore conclude the proof of Theorem 3.5.
Observe that to prove item $(5)$ it is enough to show the existence of $\Sigma_{s}^{\prime}\subset\Sigma_{s}$ such that:
  • $\Sigma_{s}^{\prime}$ is an irreducible topological Markov shift,
  • $\mathrm{proj}_{\Sigma}(\Sigma_{s}^{\prime})=\Sigma$, and
  • $\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma^{\prime}_{s})\subset\Sigma_{b}$ is a subshift of finite type.
Since $\mathcal{X}$ is a Borel homoclinic class, recall that $\Sigma$ is an irreducible topological Markov shift. Therefore, there exists $\underline{y}\in\Sigma$ such that $\Sigma=\overline{\{\sigma^{i}\left(\underline{y}\right)\}_{i\in\mathbb{Z}}}.$ Take $\underline{x}\in\Sigma_{s}$ such that $\mathrm{proj}_{\Sigma_{s}}(\underline{x})=\underline{y}$. Let $\mathscr{G}^{\prime}_{s}\subset\mathscr{G}_{s}$ the maximal irreducible component of $\mathscr{G}_{s}$ containing $x_{0}$ (the 0-th element of $\underline{x}$) and define
$\Sigma^{\prime}_{s}:=\Sigma(\mathscr{G}^{\prime}_{s})\subset\Sigma(\mathscr{G}_{s})=\Sigma_{s}.$
Observe that $\underline{x}\in\Sigma_{s}^{\prime}$ and therefore $\sigma^{i}_{s}(\underline{x})\in\Sigma_{s}^{\prime}$ for each $i\in\mathbb{Z}.$ It follows that
$\Sigma$ $=\overline{\{\sigma^{i}(\underline{y})\}_{i\in\mathbb{Z}}}=\overline{\{\sigma^{i}\circ\mathrm{proj}_{\Sigma}(\underline{x})\}_{i\in\mathbb{Z}}}=\overline{\{\mathrm{proj}_{\Sigma}\circ\sigma^{i}_{s}(\underline{x})\}_{i\in\mathbb{Z}}}=\mathrm{proj}_{\Sigma}\left(\overline{\{\sigma^{i}_{s}(\underline{x})\}_{i\in\mathbb{Z}}}\right)$
$=\mathrm{proj}_{\Sigma}(\Sigma_{s}^{\prime})\subset\mathrm{proj}_{\Sigma}(\Sigma_{s})=\Sigma.$
In this way $\Sigma=\mathrm{proj}_{\Sigma}(\Sigma_{s}^{\prime})$. As in (7), we have that $\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s}^{\prime})$ is an subshift of finite type. Moreover, since $\Sigma^{\prime}_{s}$ is irreducible, it is clear that $\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s}^{\prime})$ will also be irreducible, which concludes the proof of the theorem.
4. Measures of Relative Maximal Entropy on Topological Markov Shifts
This section is devoted to proving Theorem 4.4 and Corollary 4.5. Since the results are formulated in an abstract shift setting, we adopt notation consistent with the literature on relative entropy for shifts.
Definition 4.1.
Given
  1. (1)
    $(X,\sigma_{X})$ and $(Y,\sigma_{Y})$ shift spaces;
  2. (2)
    $\pi:X\to Y$ a surjective $1$-block factor map;
  3. (3)
    a $\sigma_{Y}$-invariant measure $\nu$ on $Y$;
  4. (4)
    two $\sigma_{X}$-invariant measures $\mu_{1}$ and $\mu_{2}$ on $X$ with $\pi_{*}\mu_{1}=\pi_{*}\mu_{2}=\nu$;
we define the $\nu$-relatively independent joining of $\mu_{1}$ and $\mu_{2}$ as the unique measure on $X\times X$ satisfying
$\mu_{1}\otimes_{\nu}\mu_{2}(A_{1}\times A_{2})=\int_{Y}\mathbb{E}_{\mu_{1}}[\mathbbm{1}_{A_{1}}\mid\pi^{-1}\mathcal{B}(Y)]\circ\pi^{-1}(y)\cdot\mathbb{E}_{\mu_{2}}[\mathbbm{1}_{A_{2}}\mid\pi^{-1}\mathcal{B}(Y)]\circ\pi^{-1}(y)\,\nu(\mathrm{d}y).$
for each $A_{1},A_{2}\in\mathcal{B}(X)$.
In the case that $X$ is a countable topological Markov shift and $\mu$ is a $\sigma_{X}$-invariant probability measure, we denote $h_{\mathrm{top}}(\sigma_{X})$ as its Gurevich topological entropy and $h_{\mu}(\sigma_{X})$ its metric entropy, for details see [41] (see also, [23, 22]).
Definition 4.2 ([39]).
We say that $\mu_{1}$ and $\mu_{2}$ are $\nu$-relative orthogonal if
$\mu_{1}\otimes_{\nu}\mu_{2}\left(\{(u,v)\in X\times X;u_{0}=v_{0}\}\right)=0$
Notation 4.3.
Given a topological Markov shift $(X,\sigma)$, we denote as $\mathcal{A}_{X}$ its alphabet. Also, given $\underline{x}\in X$ we denote as $x_{i}$ its $i$-th coordinate. Let $a\in A_{X}$, we define the $1$-cylinder at coordinate $k$ as $a]_{k}:=\{\underline{x}\in X;x_{k}=a\}.$
The purpose of this section is to show the following theorem.
Theorem 4.4.
Let $(X,\sigma_{X})$ be an irreducible countable topological Markov shift with finite Gurevich entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a surjective $1$-block factor map. Then, if $\mu_{1}$ and $\mu_{2}$ are ergodic measures of $\nu$-maximal relative entropy then $\mu_{1}$ and $\mu_{2}$ are $\nu$-relative orthogonal.
Before proving Theorem 4.4 we highlight the following useful corollary:
Corollary 4.5.
Let $(X,\sigma_{X})$ be an irreducible countable topological Markov shift with finite Gurevich entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a surjective $1$-block factor map. Then, $(X,\sigma_{X})$ can only admit countably many ergodic measures of $\nu$-maximal relative entropy.
Proof.
Assume that $(X,\sigma_{X})$ admits uncountably many $\nu$-relative measures of maximal entropy $\{\mu_{i}\}_{i\in\mathcal{I}}$. Since $\mathcal{I}$ is uncountable while the alphabet $\mathcal{A}_{X}$ is countable, there exist $a\in\mathcal{A}_{X}$ and an uncountable subset $\mathcal{I}_{0}\subset\mathcal{I}$ such that
$\mu_{i}([a]_{0})>0\text{ for every }i\in\mathcal{I}_{0}.$
For each $i\in\mathcal{I}_{0}$ we have that
$\mu_{i}([a]_{0})=\int_{Y}\mathbb{E}_{\mu_{i}}\left[\mathbbm{1}_{[a]_{0}}\mid\pi^{-1}\mathcal{B}(Y)\right]\circ\pi^{-1}(y)\,\nu(\mathrm{d}y)>0.$
Let $g_{i}(y):=\mathbb{E}_{\mu_{i}}\left[\mathbbm{1}_{[a]_{0}}\mid\pi^{-1}\mathcal{B}(Y)\right](\pi^{-1}(y))$; then with $A_{i}:=\{y\in Y;\,g_{i}(y)>0\}$ we have $\nu(A_{i})>0$.
By Theorem 4.4, for any distinct $i_{1},i_{2}\in\mathcal{I}_{0}$ the $\nu$-relatively independent joining $\mu_{i_{1}}\otimes_{\nu}\mu_{i_{2}}$ gives zero mass to the coincidence event $\{(u,v)\in X\times X;\,u_{0}=v_{0}\}$; in particular,
$\mu_{i_{1}}\otimes_{\nu}\mu_{i_{2}}\bigl(\{(u,v);\,u_{0}=v_{0}=a\}\bigr)=0.$
From the definition of $\mu_{i_{1}}\otimes_{\nu}\mu_{i_{2}}$ given in Definition 4.1 we obtain that $\nu(A_{i_{1}}\cap A_{i_{2}})=0.$ Hence the family $\{A_{i}\}_{i\in\mathcal{I}_{0}}$ is pairwise disjoint modulo $\nu$-null sets.
Moreover,
$\mathcal{I}_{0}=\bigcup_{n\in\mathbb{N}}\Bigl\{i\in\mathcal{I}_{0};\,\nu(A_{i})>\tfrac{1}{n}\Bigr\}.$
For some $n$ the set $\mathcal{I}_{1}:=\{i\in\mathcal{I}_{0};\,\nu(A_{i})>\tfrac{1}{n}\}$ is uncountable. But in a probability space there can be at most $n$ pairwise disjoint measurable sets of measure >1/n>1/n, a contradiction. Therefore $(X,\sigma_{X})$ cannot admit uncountably many $\nu$-relative measures of maximal entropy.
From now on, we fix $X$ as an irreducible countable-state topological Markov shift, $Y$ as an irreducible subshift of finite type, and $\pi:X\to Y$ a continuous surjective $1$-block factor map.
In the following, we will construct a subshift of finite type $X_{n}$ which, in some sense, approximates $X$. Since $\pi$ is a $1$-block factor map we have that $\gamma:=\{\pi^{-1}([i]_{0})\in\mathcal{B}(X);i\in\mathcal{A}_{Y}\}$ is a partition of $X$ such that each $\gamma$ is a union of $1$-cylinders at coordinate 0. For each $n\in\mathbb{N}$ we define the partition of $X$ as
$\alpha_{n}:=\left\{[a_{1}],[a_{2}],\ldots,[a_{n}],X\setminus\left(\bigcup_{i=1}^{n}[a_{i}]\right)\right\}\bigvee\gamma,$
where $\{a_{i};i\in\mathbb{N}\}$ is a fixed enumeration of $\mathcal{A}_{X}$. Observe that each element of $\alpha_{n}$ is the union of $1$-cylinders at coordinate 0.
Using the partition $\alpha_{n}$, we define a subshift of finite type $X_{n}$ as follows. Let $\mathscr{G}_{n}=(\mathscr{V}_{n},\mathscr{E}_{n})$ be the directed graph with vertex set $\mathscr{V}_{n}:=\alpha_{n}$ and edge set
$\mathscr{E}_{n}:=\bigl\{A\to A^{\prime};\,A,A^{\prime}\in\alpha_{n}\ \text{and }A\cap\sigma_{X}^{-1}(A^{\prime})\neq\emptyset\bigr\}.$
We then set $X_{n}:=\Sigma(\mathscr{G}_{n})$.
Define $\mathrm{proj}_{X_{n}}:X\to X_{n}$ as the canonical $1$-block code associated with $\alpha_{n}$:
$(\mathrm{proj}_{X_{n}}(\underline{x}))_{k}=A\ \text{whenever}\ \sigma_{X}^{k}(x)\in A,\ A\in\alpha_{n},\ k\in\mathbb{Z}.$
Also, since the partition $\alpha_{n}$ refines the partition $\gamma$, for each $A\in\alpha_{n}$ there exists a unique $i(A)\in\{[i_{0}];i_{0}\in\mathcal{A}_{Y}\}$ such that $\pi(A)\subset[i(A)]_{0}\subset Y.$ This labels each atom $A\in\alpha_{n}$ by $i(A)$ and induces a $1$-block factor map $\pi_{n}:X_{n}\to Y$ (induced by $\pi$ defined by
$\pi_{n}:(A_{k})_{k\in\mathbb{N}}\in X_{n}\mapsto(i(A_{k}))_{k\in\mathbb{N}}\in Y.$
Observe that for each $n\in\mathbb{N}$ we have that $\pi=\pi_{n}\circ\mathrm{proj}_{X_{n}}.$
Given an invariant measure $\mu$ of $(X,\sigma)$ such that $\pi_{*}\mu=\nu$, observe that $\mu_{n}:=(\mathrm{proj}_{X_{n}})_{*}\mu$ satisfies $(\pi_{n})_{*}\mu_{n}=\nu$. From the Kolmogorov-Sinai theorem (see [46, Theorem 4.9]) we have that
[…]
where $\beta_{n}$ is the partition of $X_{n}$ 11-cylinders at coordinate 0. Since $\mathcal{B}(X)=\mathcal{B}\left(\bigcup_{n}\sigma(\alpha_{n})\right)$ from [46, Theorem 4.14]
$h_{\mu}(\sigma_{X})=\lim_{n\to\infty}h_{\mu_{n}}(\sigma_{X_{n}}).$
4.1. Petersen-Quas-Shin construction
Fix $n\in\mathbb{N}$. Let $\mu_{n}^{1},\mu_{n}^{2}$ be two distinct ergodic $\sigma_{X_{n}}$-invariant measures on $X_{n}$ with $(\pi_{n})_{*}\mu_{n}^{1}=(\pi_{n})_{*}\mu_{n}^{2}=\nu,$ and assume $\mu_{n}^{1}$ and $\mu_{n}^{2}$ are not $\nu$-relatively orthogonal (Definition 4.2). Following [39], there exists an invariant measure $\mu_{n}^{3}$ on $X_{n}$ with $(\pi_{n})_{*}\mu_{n}^{3}=\nu$ and
$h_{\mu_{n}^{3}}(\sigma_{X_{n}})>\tfrac{1}{2}\left(h_{\mu_{n}^{1}}(\sigma_{X_{n}}\right)+h_{\mu_{n}^{2}}\left(\sigma_{X_{n}})\right).$
We sketch the construction of $\mu_{n}^{3}$. Diagram 2 summarises all the factor maps used in this construction in this section. Let $Z_{n}:=X_{n}\times X_{n}\times\{0,1\}^{\mathbb{Z}}$ and define the map $T:=\sigma_{X_{n}}\times\sigma_{X_{n}}\times\sigma_{\mathrm{Ber}}$, where $\sigma_{\mathrm{Ber}}$ is the shift on $\{0,1\}^{\mathbb{Z}}$. Let $\zeta$ be the (1/2,1/2)(1/2,1/2)–Bernoulli measure on $\{0,1\}^{\mathbb{Z}}$. Also, set $\widehat{\mu}_{n}:=\mu_{n}^{1}\otimes_{\nu}\mu_{n}^{2}$ and $\eta_{n}=\widehat{\mu}_{n}\otimes\zeta$. Observe that $\eta_{n}$ is TT–invariant.
Define three maps $\mathrm{proj}_{X_{n}}^{1},\mathrm{proj}_{X_{n}}^{2},\mathrm{proj}_{X_{n}}^{3}:Z_{n}\to X_{n}$ by
  1. (i)
    $\mathrm{proj}_{X_{n}}^{1}(u,v,r):=u$,
  2. (ii)
    $\mathrm{proj}_{X_{n}}^{2}(u,v,r):=v$,
  3. (iii)
    for $\mathrm{proj}_{X_{n}}^{3}$, first set, for each $k\in\mathbb{Z}$,
    $N_{k}(u,v):=\sup\{m<k:u_{m}=v_{m}\},\,\text{with the convention }\sup\emptyset=-\infty.$
    Let $r_{-\infty}:\{0,1\}^{\mathbb{Z}}\to\{0,1\}$ be a random variable taking the values 0 or $1$ according to the law $\zeta$. Define
    $\bigl(\mathrm{proj}_{X_{n}}^{3}(u,v,r)\bigr)_{k}:=\begin{cases}u_{k},&\text{if }r_{N_{k}(u,v)}=0,\\[2.0pt] v_{k},&\text{if }r_{N_{k}(u,v)}=1.\end{cases}$
By construction, $\mathrm{proj}_{X_{n}}^{3}(u,v,r)\in X_{n}$: between successive coincidence times of $u$ and $v$ we copy a block from either $u$ or $v$, and since $X_{n}$ is a topological Markov shift, switching only at coincidence times preserves admissibility. From the definition of $\widehat{\mu}_{n}$, we have that
$\pi_{n}(u)=\pi_{n}(v)\mbox{ for }\widehat{\mu}_{n}\text{-almost every $(u,v)\in X_{n}\times X_{n}$}.$
Since $\pi_{n}$ is 1-block map we obtain that
$\pi_{n}\circ\mathrm{proj}_{X_{n}}^{3}(z)=\pi_{n}\circ\mathrm{proj}_{X_{n}}^{2}(z)=\pi_{n}\circ\mathrm{proj}_{X_{n}}^{1}(z)\ \text{for }\eta_{n}\text{-almost every}\ z\in Z_{n}.$ (8)
Observe that by construction $\mu_{n}^{1}=(\mathrm{proj}_{X_{n}}^{1})_{*}\eta_{n}$ and $\mu_{n}^{2}=(\mathrm{proj}_{X_{n}}^{2})_{*}\eta_{n}$. We therefore define
$\mu_{n}^{3}=(\mathrm{proj}_{X_{n}}^{3})_{*}\eta_{n},$
Since $\mathrm{proj}_{X_{n}}^{3}:Z_{n}\to X_{n}$ is a $1$-block factor map, $\mu_{n}^{3}$ is $\sigma_{X_{n}}$-invariant.
Now, we quote a result from [39] which is contained in the proof of Theorem 1 (see [39, Equation (45)]).
Theorem 4.6 ([39]).
Under the above assumptions and notations, for each $n\in\mathbb{N}$
$h_{\mu_{n}^{3}}(\sigma_{X_{n}})-h_{\nu}(\sigma_{Y})\geq\frac{h_{\mu_{n}^{1}}(\sigma_{X_{n}})+h_{\mu_{n}^{2}}(\sigma_{X_{n}})}{2}-h_{\nu}(\sigma_{Y})+\int_{S_{n}}\Xi_{n}(z)\,\eta_{n}(\mathrm{d}z),$
where
$S_{n}:=\left\{z\in Z_{n};\,(\mathrm{proj}_{X_{n}}^{1}(z))_{-1}=(\mathrm{proj}_{X_{n}}^{2}(z))_{-1}\,\right\}.$
The integrand $\Xi_{n}(z)$ is given by
$\Xi_{n}(z)=\sum_{j\in\mathcal{A}_{X_{n}}}\psi\left(\frac{A_{n}^{j}(z)+B_{n}^{j}(z)}{2}\right)-\frac{1}{2}\sum_{j\in\mathcal{A}_{X_{n}}}\psi\big(A_{n}^{j}(z)\big)-\frac{1}{2}\sum_{j\in\mathcal{A}_{X_{n}}}\psi\big(B_{n}^{j}(z)\big),$
with $\psi(t):=-\,t\log t$, and
$A_{n}^{j}(z):=\mathbb{E}_{\eta_{n}}\big(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\,\big|\,\mathcal{B}_{1,n}^{-}\vee\mathcal{B}_{0,n}\big)(z),\ B_{n}^{j}(z):=\mathbb{E}_{\eta_{n}}\big(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{2}\,\big|\,\mathcal{B}_{2,n}^{-}\vee\mathcal{B}_{0,n}\big)(z),$
where
$\mathcal{B}_{1,n}^{-}:=\sigma\left(\,\{z\in Z_{n};\ (\mathrm{proj}_{X_{n}}^{1}(z))_{-m}=a\,\};\,a\in\mathcal{A}_{X_{n}},\ m\in\mathbb{N}\right),$
$\mathcal{B}_{2,n}^{-}:=\sigma\left(\,\{z\in Z_{n};\ (\mathrm{proj}_{X_{n}}^{2}(z))_{-m}=a\,\};\,a\in\mathcal{A}_{X_{n}},\ m\in\mathbb{N}\right),$
$\mathcal{B}_{0,n}:=(\pi_{n}\circ\mathrm{proj}_{X_{n}}^{1})^{-1}(\mathcal{B}(Y)).$ (9)
Remark 4.7.
Observe that by (8) one has
$(\pi_{n}\circ\mathrm{proj}_{X_{n}}^{1})^{-1}(\mathcal{B}(Y))=(\pi_{n}\circ\mathrm{proj}_{X_{n}}^{2})^{-1}(\mathcal{B}(Y))=(\pi_{n}\circ\mathrm{proj}_{X_{n}}^{3})^{-1}(\mathcal{B}(Y))\ (\mathrm{mod}\ \eta_{n}).$
Thus, we may alternatively define $\mathcal{B}_{0,n}:=(\pi_{n}\circ\mathrm{proj}_{X_{n}}^{i})^{-1}(\mathcal{B}(Y))$ in (9) using any $i\in\{1,2,3\}$, and the conclusions of Theorem 4.6 hold.
[…]
Diagram 2: Diagram of the maps $p_{X}^{i},\mathrm{proj}_{Z_{n}}$, $\mathrm{proj}_{X_{n}}$, $\pi$, $\pi_{n}$, $\mathrm{proj}_{X}^{i}$ and $\mathrm{proj}^{i}_{X_{n}}$ with $i\in\{1,2,3\}$. The right diagram commutes, and all maps are surjective semi-conjugacies between the shift maps on the respective spaces.
Below we quote another result from [39]. In [39], the statement is proved under the assumption that $X$ is a subshift of finite type, but the same proof works assuming that $X$ is a countable-state irreducible Markov shift and $Y$ is a subshift of finite type.
Lemma 4.8 ([39, Lemma 3]).
Consider a two-sided countable Markov shift $X\subset\mathcal{A}_{X}^{\mathbb{Z}}$ with shift $\sigma_{X}$, a subshift $Y$ with shift $\sigma_{Y}$, and a $1$-block factor map $\pi:X\to Y$. Fix an ergodic $\sigma_{Y}$-invariant probability $\nu$ on $Y$ and choose ergodic $\mu_{1},\mu_{2}\in\pi^{-1}\{\nu\}$. Denote by $\widehat{\mu}:=\mu_{1}\otimes_{\nu}\mu_{2}$ their relatively independent joining on $X\times X$, and by $p_{X}^{1},p_{X}^{2}:X\times X\to X$ the coordinate projections.
Put $\mathcal{F}^{-}:=\sigma(x_{k};\,k<0)$ for the past $\sigma$-algebra on $X$, and set $\mathcal{F}^{-}_{i}:=(p^{i}_{X})^{-1}(\mathcal{F}^{-})$, $i\in\{1,2\}$ on $X\times X$. Write $\mathcal{Q}_{Y}:=\sigma([y_{0}]_{0};\,y_{0}\in\mathcal{A}_{Y})$ and define
[…]
Assume that $\widetilde{S}:=\{(u,v)\in X\times X;\ u_{-1}=v_{-1}\}$ has positive $\widehat{\mu}-$measure, and
$\mathbb{E}_{\widehat{\mu}}\bigl(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X}^{1}\,\bigm|\,\mathcal{F}^{-}_{1}\vee\mathcal{F}_{0}\bigr)=\mathbb{E}_{\widehat{\mu}}\bigl(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X}^{2}\,\bigm|\,\mathcal{F}^{-}_{2}\vee\mathcal{F}_{0}\bigr)\ \widehat{\mu}\text{-a.e. on }\widetilde{S}$,
for every $j\in\mathcal{A}_{X}$, then $\mu_{1}=\mu_{2}$.
Proof of Theorem 4.4.
We assume the notations of Theorem 4.6 and Lemma 4.8. Assume, for contradiction, that the theorem is false. Then there exist two ergodic measures of $\nu$-maximal relative entropy, $\mu^{1}$ and $\mu^{2}$, such that
$\mu^{1}\otimes_{\nu}\mu^{2}\left(\{(u,v)\in X\times X;\,u_{0}=v_{0}\}\right)>0.$
Define $\mu^{3}=(\mathrm{proj}^{3}_{X_{n}})_{*}\eta$ by the construction at the beginning of Section 4.1 (we keep the same notation, omitting the index $n$). We will show that
$h_{\mu^{3}}(X)>h_{\mu^{1}}(\sigma_{X})=h_{\mu^{2}}(\sigma_{X})=h_{\mathrm{top}}(\sigma_{X}\mid\nu).$
which yields a contradiction. Note that we cannot directly apply the result of Lemma 4.8 since $X$ is not a subshift of finite type.
From the discussion following the proof of Corollary 4.5,
$h_{\mu^{i}}(X)=\lim_{n\to\infty}h_{\mu^{i}_{n}}(X_{n})\ \text{for }i\in\{1,2,3\},\ \text{where}\ \mu_{n}^{i}:=(\pi_{n})_{*}\mu^{i}.$
Since $\mu^{1}\otimes_{\nu}\mu^{2}$ is invariant under $\sigma_{X}\times\sigma_{X}$,
$\mu^{1}\otimes_{\nu}\mu^{2}\left(\{(u,v)\in X\times X;\,u_{-1}=v_{-1}\}\right)>0.$
By Lemma 4.8 there exists $j_{0}\in\mathcal{A}_{X}$ such that, on a $\mu^{1}\otimes_{\nu}\mu^{2}$-positive subset of $\{(u,v)\in X\times X:u_{-1}=v_{-1}\}$,
$\mathbb{E}_{\mu^{1}\otimes_{\nu}\mu^{2}}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ p_{X}^{1}\,\bigm|\,\mathcal{F}^{-}_{1}\vee\mathcal{B}_{0}\bigr)\neq\mathbb{E}_{\mu^{1}\otimes_{\nu}\mu^{2}}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ p_{X}^{2}\,\bigm|\,\mathcal{F}^{-}_{2}\vee\mathcal{B}_{0}\bigr).$
Hence
$\mathbb{E}_{\eta}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ\mathrm{proj}_{X}^{1}\,\bigm|\,\mathcal{B}^{-}_{1}\vee\mathcal{B}_{0}\bigr)\neq\mathbb{E}_{\eta}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ\mathrm{proj}_{X}^{2}\,\bigm|\,\mathcal{B}^{-}_{2}\vee\mathcal{B}_{0}\bigr)$ (10)
on an $\eta$-positive measure set contained in $S:=\{\,z\in Z;\ (\mathrm{proj}_{X}^{1}z)_{-1}=(\mathrm{proj}_{X}^{2}z)_{-1}\,\}.$
Let $\mathrm{proj}_{Z_{n}}:Z\to Z_{n}$ be the projection
$\mathrm{proj}_{Z_{n}}:Z$ $\to Z_{n}$
$z=(x,y,r)$ $\mapsto(\mathrm{proj}_{X_{n}}(x),\mathrm{proj}_{X_{n}}(y),r),$
and observe that
$\mathrm{proj}_{X_{n}}^{i}\circ\mathrm{proj}_{Z_{n}}=\mathrm{proj}_{X_{n}}\circ\mathrm{proj}_{X}^{i}\ \text{for any }i\in\{1,2,3\}.$
For each $i\in\{1,2\}$ the sequence of $\sigma$-algebras
$(\mathrm{proj}_{Z_{n}})^{-1}\bigl(\mathcal{B}_{i,n}^{-}\vee\mathcal{B}_{0,n}\bigr),\qquad n\geq 1,$
is increasing and
$\sigma\left(\bigcup_{n\geq 1}(\mathrm{proj}_{Z_{n}})^{-1}\bigl(\mathcal{B}_{i,n}^{-}\vee\mathcal{B}_{0,n}\bigr)\right)=\mathcal{B}_{i}^{-}\vee\mathcal{B}_{0}.$
By Doob’s martingale convergence theorem,
$\mathbb{E}_{\eta}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ\mathrm{proj}_{X}^{i}\,\bigm|\,\mathcal{B}^{-}_{i}\vee\mathcal{B}_{0}\bigr)=\lim_{n\to\infty}\mathbb{E}_{\eta}\left(\mathbbm{1}_{[j_{0}]_{0}}\circ\mathrm{proj}_{X}^{i}\,\bigm|\,(\mathrm{proj}_{Z_{n}})^{-1}(\mathcal{B}_{i,n}^{-}\vee\mathcal{B}_{0,n})\right)$ (11)
in $L^{1}(Z,\eta)$ and $\eta$-a.s. Moreover,
$\mathrm{proj}_{Z_{n}}(S)\subset S_{n}:=\{\,z\in Z_{n};\ (\mathrm{proj}_{X_{n}}^{1}(z))_{-1}=(\mathrm{proj}_{X_{n}}^{2}(z))_{-1}\,\}.$
With $\psi(t):=-t\log t$,
$\int_{S_{n}}\psi\left(\frac{A_{n}^{j_{0}}+B_{n}^{j_{0}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}})-\frac{1}{2}\psi(B_{n}^{j_{0}})\,\mathrm{d}\eta_{n}$
$=\int_{{\mathrm{proj}^{-1}_{Z_{n}}}(S_{n})}\psi\left(\frac{A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}+B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}})-\frac{1}{2}\psi(B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}})\,\mathrm{d}\eta$
$\geq\int_{S}\psi\left(\frac{A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}+B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}})-\frac{1}{2}\psi(B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}})\,\mathrm{d}\eta,$ (12)
where $A_{n}^{j},B_{n}^{j}$ are as in Theorem 4.6. Observe that
$A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}$ $=\mathbb{E}_{\eta_{n}}\big(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\,\big|\,\mathcal{B}_{1,n}^{-}\vee\mathcal{B}_{0,n}\big)\circ\mathrm{proj}_{Z_{n}}$
$=\mathbb{E}_{\eta}\left(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X}^{1}\,\big|\,(\mathrm{proj}_{Z_{n}})^{-1}\left(\mathcal{B}_{1,n}^{-}\vee\mathcal{B}_{0,n}\right)\right)$ (13)
Similarly, we obtain that
$B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}=\mathbb{E}_{\eta}\left(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X}^{2}\,\big|\,(\mathrm{proj}_{Z_{n}})^{-1}\left(\mathcal{B}_{2,n}^{-}\vee\mathcal{B}_{0,n}\right)\right).$ (14)
Combining (12) with (13), (14), (10), and (11) yields
$\lim_{n\to\infty}\int_{S_{n}}\left[\psi\left(\frac{A_{n}^{j_{0}}+B_{n}^{j_{0}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}})-\frac{1}{2}\psi(B_{n}^{j_{0}})\right]\mathrm{d}\eta_{n}>0.$ (15)
Finally, Theorem 4.6 together with (15) gives
$h_{\mu^{3}}(\sigma_{X})$ $=\lim_{n\to\infty}h_{\mu^{3}_{n}}(\sigma_{X_{n}})$
$\geq\lim_{n\to\infty}\left[\frac{h_{\mu^{1}_{n}}(\sigma_{X_{n}})+h_{\mu^{2}_{n}}(\sigma_{X_{n}})}{2}+\int_{S_{n}}\Xi_{n}(z)\,\eta_{n}(\mathrm{d}z)\right]$
$\geq\lim_{n\to\infty}\left[\frac{h_{\mu^{1}_{n}}(\sigma_{X_{n}})+h_{\mu^{2}_{n}}(\sigma_{X_{n}})}{2}+\int_{S_{n}}\psi\left(\frac{A_{n}^{j_{0}}+B_{n}^{j_{0}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}})-\frac{1}{2}\psi(B_{n}^{j_{0}})\mathrm{d}\eta_{n}\right]$
$>\frac{h_{\mu^{1}}(\sigma_{X})+h_{\mu^{2}}(\sigma_{X})}{2}=h_{\mu^{1}}(\sigma_{X}).$
This contradicts $h_{\mu^{1}}(\sigma_{X}\mid\mathbb{\nu})=h_{\mathrm{top}}(\sigma_{X}\mid\nu)$, since $\pi_{*}\mu^{3}=\nu$.
5. Proofs of Theorem A, Corollary B and Proposition 1.6
Before proving Theorem A, we quote two results from the literature and prove a lemma. The first is a consequence of [49, Theorems 4.4 and 5.1] and the second is a result proved in [40].
Lemma 5.1 ([49, Lemma 3.1, Theorems 4.4 and 5.1]).
Let $(\Sigma^{\prime}_{b},\sigma_{b})$ be a subshift of finite type, $\Lambda$ be topological space, $\theta:\Lambda\to\Lambda$ a homeomorphism, $\pi:\Sigma\to\Lambda$ a surjective finite-to-one Hölder continuous map. Then,
  • (i)
    for each ergodic $\mathbb{P}\in\mathcal{M}_{e}(\theta)$ there exists $\widehat{\mathbb{P}}\in\mathcal{M}_{e}(\sigma)$ such that $(\pi)_{*}\widehat{\mathbb{P}}$$\mathbb{P}$;
  • (ii)
    there exists only finitely many ergodic probability measures $\widehat{\mathbb{P}}_{1},\ldots,\widehat{\mathbb{P}}_{k}\in\mathcal{M}_{e}(\theta)$, such that $(\pi)_{*}\widehat{\mathbb{P}}_{i}=\mathbb{P}$; and
  • (iii)
    there exists $N_{\mathbb{P}}\in\mathbb{N}$ such that $\pi^{-1}(\omega)=N_{\mathbb{P}}$ for $\mathbb{P}$-almost every $\omega\in\Lambda$.
We note that [40, Proposition 13.2] treats a slightly different setting, but the proof applies in our context because it uses only the finite-to-one coding properties (see also [9, Thm. 0.2] and [18, Thm. 10.5]).
Lemma 5.2 ([40, Proposition 13.2 and Step 2 of its proof]).
Let $\Theta$ be as in Theorem A and fix $\chi>0$. Consider $\Sigma_{s}$ and $\pi_{s}:\Sigma_{s}\to\Omega\times M$ as in Theorem 3.5. Then, for each $\chi$-hyperbolic $\mu\in\mathcal{M}_{e}(\Theta)$
  1. (i)
    there exists $\widehat{\mu}\in M_{e}(\sigma_{s})$ such that $(\pi_{s})_{*}\widehat{\mu}=\mu$;
  2. (ii)
    $h_{\widehat{\mu}}(\sigma_{s})=h_{\mu}(\Theta);$ and
  3. (iii)
    there exists $N_{\mu}$ such that $\pi_{s}^{-1}(\omega,x)=N_{\mu}$ for $\mu$-almost every $(\omega,x)\in\Omega\times M$.
Lemma 5.3.
Let $\Theta$ and $\mathbb{P}$ be as in Theorem A. Define
$\mathcal{M}_{\mathrm{max}}(\Theta\mid\mathbb{P}):=\left\{\mu\in\mathcal{M}_{e}(\Theta)\middle|\;\begin{aligned} &(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}\ \text{and }\mu\ \text{is a hyperbolic}\\ &\text{measure of }\mathbb{P}\text{-maximal relative entropy}\end{aligned}\ \right\},$
If, for each $\chi>0$ and every homoclinic class $\mathcal{X}$, the set
[…]
is countable, then $\mathcal{M}_{\mathrm{max}}(\Theta\mid\mathbb{P})$ is countable.
Proof.
Define $\mathcal{M}_{\mathrm{max}}^{\chi}(\Theta\mid\mathbb{P})$ as in (16) but changing $\mathcal{X}$ to $\Omega\times M$. Observe that
$\mathcal{M}_{\mathrm{max}}(\Theta\mid\mathbb{P})=\bigcup_{\chi>0}\mathcal{M}_{\mathrm{max}}^{\chi}(\Theta\mid\mathbb{P})=\bigcup_{n\in\mathbb{N}}\mathcal{M}_{\mathrm{max}}^{\frac{1}{n}}(\Theta\mid\mathbb{P}).$
Thus it suffices to show that each $\mathcal{M}_{\mathrm{max}}^{\chi}(\Theta\mid\mathbb{P})$ is countable.
Fix $\chi>0$. Let $\mu\in\mathcal{M}_{e}(\Theta)$ be $\chi$-hyperbolic. By Theorem 2.5, $\mu\big(\mathrm{NUH}^{\prime}(\Theta)\big)=1.$ From Proposition 2.9, there exists a countable family of Borel homoclinic classes $\{\mathcal{X}_{i}\}_{i\in\mathcal{I}}$ forming a pairwise disjoint, $\Theta$-invariant partition of $\mathrm{NUH}^{\prime}(\Theta)$. Since $\mu$ is ergodic, there is a unique $i_{0}\in\mathcal{I}$ such that $\mu(\mathcal{X}_{i_{0}})=1.$ Consequently,
$\mathcal{M}_{\mathrm{max}}^{\chi}(\Theta\mid\mathbb{P})=\bigcup\left\{\mathcal{M}_{\mathrm{max}}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})\ ;\ \mathcal{X}\text{ is a Borel homoclinic class}\right\}=\bigcup_{i\in\mathcal{I}}\mathcal{M}_{\mathrm{max}}^{\mathcal{X}_{i},\chi}(\Theta\mid\mathbb{P}).$
By hypothesis, each $\mathcal{M}_{\mathrm{max}}^{\mathcal{X}_{i},\chi}(\Theta\mid\mathbb{P})$ is countable, and $\mathcal{I}$ is countable; hence $\mathcal{M}_{\mathrm{max}}^{\chi}(\Theta\mid\mathbb{P})$ is countable. Taking the countable union over $n\geq 1$ gives that $\mathcal{M}_{\mathrm{max}}(\Theta\mid\mathbb{P})$ is countable.
Now we can prove Theorem A.
Proof of Theorem A.
Let $\mathbb{P}$ be a $\theta$-invariant ergodic measure. By Lemma 5.3, it suffices to show that $\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})$ is countable for each homoclinic class $\mathcal{X}$ and each $\chi>0$.
Fix $\chi>0$ and a Borel homoclinic class $\mathcal{X}$. Let $\Sigma_{s}$ be an irreducible countable-state topological Markov shift, $\Sigma_{b}$ a subshift of finite type, and let $\pi_{s}:\Sigma_{s}\to\mathcal{X}$ and $\pi_{b}:\Sigma_{b}\to\Omega$ be the Hölder maps given by Theorem 3.5. Recall that:
  1. (1)
    $\pi_{s}\circ\sigma_{s}=\Theta\circ\pi_{s}$ and $\pi_{b}\circ\sigma_{b}=\theta\circ\pi_{b}$, where $\sigma_{s}$ and $\sigma_{b}$ denote the left shift maps on $\Sigma_{s}$ and $\Sigma_{b}$, respectively;
  2. (2)
    $\pi_{s}[\Sigma_{s}^{\#}]$ has full measure for every $\chi$-hyperbolic $\mu\in\mathcal{M}_{e}(\Theta)$ with $\mu(\mathcal{X})=1$; moreover $\pi_{s}:\Sigma_{s}^{\#}\to\mathcal{X}$ is finite-to-one, and $\pi_{b}:\Sigma_{b}\to NW(\theta)$ is a surjective finite-to-one map;
  3. (3)
    there exists a $1$-block map $\mathrm{proj}_{\Sigma_{b}}:\Sigma_{s}\to\Sigma_{b}$ such that
    $\pi_{b}\circ\mathrm{proj}_{\Sigma_{b}}=\mathrm{proj}_{\Omega}\circ\pi_{s},$
    and $\Sigma_{b}^{\prime}:=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})$ is an irreducible subshift of finite type.
Let $\mu\in\mathcal{M}_{e}(\Theta)$ be a hyperbolic measure of $\mathbb{P}$-maximal relative entropy. By Lemma 5.2, there exists $\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s})$ such that $(\pi_{s})_{*}\widehat{\mu}=\mu$ and $h_{\mu}(\Theta)=h_{\widehat{\mu}}(\sigma_{s}).$ Observe that
$(\pi_{b})_{*}\big[(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}\big]=(\pi_{b}\circ\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}=(\mathrm{proj}_{\Omega})_{*}(\pi_{s})_{*}\widehat{\mu}=(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}.$
Thus $\widehat{\mathbb{P}}:=(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$ is ergodic as a factor of $\widehat{\mu}$, and is a $\sigma_{b}$-invariant $\pi_{b}$-lift of $\mathbb{P}$ supported on $\Sigma_{b}^{\prime}$.
Claim. For the factor map $\mathrm{proj}_{\Sigma_{s}}:(\Sigma_{s},\sigma_{s})\to(\Sigma_{b}^{\prime},\sigma_{b})$, $\widehat{\mu}$ is a measure of maximal $^\widehat{\mathbb{P}}$-relative entropy. When the context is clear, we simply say that $\widehat{\mu}$ is a measure of maximal entropy relative to $\widehat{P}$.
[…]
$(\mathrm{proj}_{\Omega})_{*}\eta=(\mathrm{proj}_{\Omega}\circ\pi_{s})_{*}\widehat{\eta}=(\pi_{b})_{*}(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\eta}=(\pi_{b})_{*}\widehat{\mathbb{P}}$$\mathbb{P}$.
By Lemmas 5.1(iii) and 5.2(iii), $\pi_{b}$ is $N_{\mathbb{P}}$-to-one on a set of $\mathbb{P}$-full measure and $\pi_{s}$ is $N_{\eta}$-to-one on a set of $\eta$-full measure; since finite extensions preserve entropy, we have
[…]
$<h_{\widehat{\eta}}(\sigma_{s}\mid\widehat{\mathbb{P}})=h_{\widehat{\eta}}(\sigma_{s})-h_{\widehat{\mathbb{P}}}(\sigma_{b})=h_{\eta}(\Theta)-h_{\mathbb{P}}(\theta)=h_{\eta}(\Theta\mid\mathbb{P}).$
Since $\eta(\mathcal{X})=\widehat{\eta}\big(\pi_{s}^{-1}(\mathcal{X})\big)=\widehat{\eta}(\Sigma_{s})=1,$ Proposition 2.9 implies that $\eta$ is hyperbolic, which contradicts the maximality of $\mu$ for $\mathbb{P}$-relative entropy. Therefore $\widehat{\mu}$ is a measure of $^\widehat{\mathbb{P}}$-maximal relative entropy. ∎
For each $\mu\in\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})$ we have thus associated a measure $\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s})$ such that $\widehat{\mu}$ is a measure of maximal entropy relative to $\widehat{\mathbb{P}}:=(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$ and
$(\pi_{b}\circ\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$$\mathbb{P}$.
Define
$\mathcal{M}_{\max}(\Sigma_{s}\mid\mathbb{P}):=\left\{\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s})\ \middle|\ \begin{aligned} &\widehat{\mu}\ \text{has maximal entropy relative to }(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}\\ &\text{and }(\pi_{b}\circ\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$$\mathbb{P}\end{aligned}\right\}$.
Then, the assignment
$\mu\in\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})\mapsto\widehat{\mu}\in\mathcal{M}_{\max}(\Sigma_{s}\mid\mathbb{P})$ (17)
is well defined and injective, since $(\pi_{s})_{*}\widehat{\mu}=\mu$.
By Lemma 5.1, there exist only finitely many ergodic $\sigma_{b}$-invariant probability measures $\widehat{\mathbb{P}}_{1},\dots,\widehat{\mathbb{P}}_{k}$ supported on $\Sigma_{b}^{\prime}$ such that $(\pi_{b})_{*}\widehat{\mathbb{P}}_{i}=\mathbb{P}$. Hence, using the injectivity of (17),
$\#\,\mathcal{M}_{\max}^{\mathcal{X},\chi}$ $(\Theta\mid\mathbb{P})\leq\#\,\mathcal{M}_{\max}(\Sigma_{s}\mid\mathbb{P})$
$=\#\left(\bigcup_{i=1}^{k}\left\{\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s});\ \widehat{\mu}\ \text{has maximal entropy relative to }\widehat{\mathbb{P}}_{i}\right\}\right).$ (18)
From Corollary 4.5, with $X=\Sigma_{s}$, $Y=\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})$, $\pi=\mathrm{proj}_{\Sigma_{b}}$ and $\nu=\widehat{\mathbb{P}}_{i}$ it follows that, for each $i\in\{1,\dots,k\}$, the set
$\left\{\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s}):\,\widehat{\mu}\ \text{has maximal entropy relative to }\widehat{\mathbb{P}}_{i}\right\}$
is at most countable. Therefore from (18) we obtain that $\mathcal{M}_{\max}^{\mathcal{X},\chi}(\Theta\mid\mathbb{P})$ is countable, which completes the proof. ∎
To prove Corollary B, we recall the following theorem for skew products, which refines Ruelle’s inequality. We state a convenient weak version of the result presented in [24] that suffices for our purposes.
Theorem 5.4 ([24, Chapter 5, Theorem 3.1.1]).
Let
$\Theta:(\omega,x)\in\Omega\times M\mapsto\left(\theta(\omega),T_{\omega}(x)\right)\in\Omega\times M$
be a $\mathcal{C}^{1+\alpha}$ skew-product diffeomorphism, and let $\mu\in\mathcal{M}_{e}(\Theta)$. Denote the base marginal by $\mathbb{P}:=(\mathrm{proj}_{\Omega})_{*}\mu$. Let $\sigma(\mu)=\{\lambda_{1},\ldots,\lambda_{n}\}$ be the Lyapunov spectrum of $\Theta$ (see Definition 2.1). Define the set of fibre Lyapunov exponents
$\sigma(\mu\mid\mathbb{P}):=\Bigl\{\lambda\in\sigma(\mu);\ \mathrm{proj}_{T_{x}M}\bigl(E^{\lambda}_{(\omega,x)}\bigr)\neq\{0\}\ \text{for }\mu\text{-a.e. }(\omega,x)\in\Omega\times M\Bigr\},$
and for each $\lambda\in\sigma(\mu\mid\mathbb{P})$ set the (fibre) multiplicity
[…]
which is $\mu$-a.s. constant by ergodicity.
Then
$h_{\mu}(\Theta\mid\mathbb{P})\leq\sum_{\begin{subarray}{c}\lambda\in\sigma(\mu\mid\mathbb{P})\\ \lambda>0\end{subarray}}\lambda\,m_{\lambda}.$
We can now prove Corollary B.
Proof of Corollary B.
By virtue of Theorem A, it is sufficient to prove that every ergodic measure of $\mathbb{P}$-relative maximal entropy of $\Theta$ is hyperbolic. In the notations of Theorem 5.4, it is clear that
$\sigma(\mu)=\sigma(\theta)\cup\sigma(\mu\mid\mathbb{P}).$
Since $\theta:\Omega\to\Omega$ is Axiom $A$ then $0\notin\sigma(\theta)$. Using that $\mathrm{dim}(M)=2$ we have the cardinality of $\sigma(\mu\mid\mathbb{P})$ is either equal to $1$ or $2$. Therefore, to prove the corollary, it is enough to show that
$\sigma(\mu\mid\mathbb{P})=\{\lambda_{-},\lambda_{+}\}$
with $\lambda_{+}<0<\lambda_{+}$. From Theorem 5.4 we have that
$0<h_{\mu}(\Theta\mid\mathbb{P})\leq\sum_{\begin{subarray}{c}\lambda\in\sigma(\mu\mid\mathbb{P})\\ \lambda>0\end{subarray}}\lambda\,m_{\lambda}.$ (19)
On the other hand, applying the same result to the skew product $\Theta^{-1}$, we obtain that
$0<h_{\mu}(\Theta\mid\mathbb{P})=h_{\mu}(\Theta^{-1}\mid\mathbb{P})\leq\sum_{\begin{subarray}{c}\widetilde{\lambda}\in\sigma^{-1}(\mu\mid\mathbb{P})\\ \widetilde{\lambda}>0\end{subarray}}\widetilde{\lambda}\,m_{\widetilde{\lambda}}$, (20)
where $\sigma^{-1}(\mu\mid\mathbb{P})$ denotes the set of the fibre Lyapunov exponents of $\Theta^{-1}$. Since the fibrewise Lyapunov spectrum of $\Theta^{-1}$ is the sign-reversal of that of $\Theta$, i.e. $\sigma^{-1}(\mu\mid\mathbb{P})=\left\{-\lambda;\ \lambda\in\sigma(\mu\mid\mathbb{P})\right\}.$ It follows from (20) that
$0<h_{\mu}(\Theta\mid\mathbb{P})\leq\sum_{\begin{subarray}{c}\widetilde{\lambda}\in\sigma^{-1}(\mu\mid\mathbb{P})\\ \widetilde{\lambda}>0\end{subarray}}\widetilde{\lambda}m_{\widetilde{\lambda}}=\sum_{\begin{subarray}{c}\lambda\in\sigma(\mu\mid\mathbb{P})\\ \lambda<0\end{subarray}}-\lambda m_{\lambda}.$ (21)
From (19) and (21) combined with $\mathrm{dim}(M)=2$, we obtain that $\sigma(\mu\mid\mathbb{P})=\{\lambda_{-},\lambda_{+}\},$} which implies that $\mu$ is a hyperbolic measure and concludes the proof of the theorem.
We conclude this section by proving Proposition 1.6.
Proof of Proposition 1.6.
Consider $G_{k}:=\{x_{1}\in\mathbb{S}^{1};\,|2-2\pi k\sin(2\pi x_{1})|\geq 9\}.$ Choose $k_{0}$ large enough so that for every fixed $k\geq k_{0}$ we can pick disjoint intervals such that
$\mathrm{length}(J_{1})=\mathrm{length}(J_{2})=\frac{1}{3},\ \text{and }J_{1},J_{2}\subset G_{k}.$
Define the strips $S_{1}:=J_{1}\times\mathbb{S}^{1}$ and $S_{2}:=J_{2}\times\mathbb{S}^{1}\subset\mathbb{T}^{2}$; see Figure 1. By construction,
{ $\{\,2x_{1}+x_{2}+k\cos(2\pi x_{1})+f(\omega);\ x_{1}\in J_{i}\,\}=\mathbb{S}^{1},\text{for }i\in\{1,2\},\ \text{every }x_{2}\in\mathbb{S}^{1}\text{ and }\omega\in\mathbb{T}^{2},$
and the map $x_{1}\in J_{i}\mapsto 2x_{1}-x_{2}+k\cos(2\pi x_{1})+f(\omega)$ winds at least three times around $\mathbb{S}^{1}$ on each $i\in\{1,2\}$, since $|2-2\pi k\sin(2\pi x_{1})|\geq 9$ and $\mathrm{length}(J_{i})=1/3$.
Figure 1. Illustration of the graph transform in Steps 1–2 for $k=4$ and $\omega_{0}=0$. The grey strips indicate $S_{1}$ (left) and $S_{2}$ (right). The orange curves are the restrictions $\gamma_{1}=\gamma|_{J_{1}}\in\mathcal{G}_{1}$ and $\gamma_{2}=\gamma|_{J_{2}}\in\mathcal{G}_{2}$ of $\gamma(t)=(t,t/2)$, while the blue curves are their images $T_{\omega_{0}}\circ\gamma_{1}$ and $T_{\omega_{0}}\circ\gamma_{2}$.
Step 1.
Define, for $i\in\{1,2\}$, the family of graphs
$\mathcal{G}_{i}:=\big\{\ \gamma:J_{i}\to\mathbb{T}^{2}\ \text{ is }\mathcal{C}^{1};\ \gamma(t)=(t,g(t))\ \text{with}\ \|g^{\prime}\|_{\infty}<1\ \big\}.$
Then, for $k$ sufficiently large, for every $\omega\in\Omega$, every $\gamma\in\mathcal{G}_{i}$, and each $j\in\{1,2\}$, there exists $\widetilde{\gamma}=\widetilde{\gamma}(\omega,\gamma,j)\in\mathcal{G}_{j}$ such that
$\widetilde{\gamma}(J_{j})\subset T_{\omega}\big(\gamma(J_{i})\big).$
Indeed, for $\gamma(t)=(t,g(t))$ we compute
$T_{\omega}\circ\gamma(t)=\big(2t-g(t)+k\cos(2\pi t)+f(\omega),\ t\big).$
For $t\in J_{i}\subset G_{k}$,
$\left|\frac{d}{dt}\big(2t-g(t)+k\cos(2\pi t)+f(\omega)\big)\right|\geq|2-2\pi k\sin(2\pi t)|-|g^{\prime}(t)|\ \geq\ 9-1\ =\ 8.$
Hence $t\mapsto 2t-g(t)+k\cos(2\pi t)+f(\omega)$ is strictly monotone on $J_{i}$ and, since $\mathrm{length}(J_{i})=1/3$, its image wraps at least twice around the horizontal circle. Let $h:\mathbb{S}^{1}\rightarrow J_{i}$ be any inverse branch of
$t\in J_{1}\mapsto 2t-g(t)+k\cos(2\pi t)+f(\omega)\in\mathbb{S}^{1}$
Then
$\{(s,h(s));\ s\in J_{j}\}\subset\{\,T_{\omega}\circ\gamma(t);\ t\in J_{i}\,\}.$
Define $\widetilde{\gamma}(s):=(s,h(s))$. By the chain rule and the identity
$h\big(2t-g(t)+k\cos(2\pi t)+f(\omega)\big)=t,$
we obtain that
$\sup_{y\in J_{j}}|h^{\prime}(y)|=\sup_{t\in J_{i}}\left|\frac{1}{2-g^{\prime}(t)-2\pi k\sin(2\pi t)}\right|\leq\frac{1}{8}<1.$
Therefore, $\widetilde{\gamma}\in\mathcal{G}_{j}$ and Step 1 follows.
Step 2.
Given a sequence $\underline{r}=(r_{\ell})_{\ell\in\mathbb{Z}}\in\{1,2\}^{\mathbb{Z}}$ and any $\omega\in\Omega$, there exists $x\in\mathbb{T}^{2}$ such that $T_{\omega}^{\ell}(x)\in S_{r_{\ell}}$, for every $\ell\in\mathbb{N}.$
Recall the definition of the family $\mathcal{G}_{\ell}$ from Step 1, with $\ell\in\{1,2\}$. Consider the graph $\gamma_{0}\colon J_{r_{0}}\to\mathbb{T}^{2}$ defined by $\gamma_{0}(t)=(t,0)$. Clearly $\gamma_{0}\in\mathcal{G}_{r_{0}}$. By Step 1, there exists $\gamma_{1}\colon J_{r_{1}}\to\mathbb{T}^{2}$ with $\gamma_{1}\in\mathcal{G}_{r_{1}}$ such that
$\gamma_{1}(J_{r_{1}})\subset T_{\omega}\big(\gamma_{0}(J_{r_{0}})\big).$
Proceeding inductively, we construct graphs $\{\gamma_{\ell}\colon J_{r_{\ell}}\to\mathbb{T}^{2}\}_{\ell\in\mathbb{N}}$ with $\gamma_{\ell}\in\mathcal{G}_{r_{\ell}}$ such that for each $\ell\geq 1$,
$\gamma_{\ell}(J_{r_{\ell}})$ $\subset T_{\theta^{\ell-1}\omega}\circ\gamma_{\ell-1}(J_{r_{\ell-1}})$
$\subset\ldots\subset T_{\theta^{\ell-1}\omega}\circ\cdots\circ T_{\theta\omega}\circ T_{\omega}\circ\gamma_{0}(J_{r_{0}})=T_{\omega}^{\ell}\circ\gamma_{0}(J_{r_{0}}).$ (22)
For each $\ell\in\mathbb{N}$ define the compact sets
$\mathcal{T}_{\ell}:=\big\{\,t\in J_{r_{0}};\ T_{\omega}^{j}\big(\gamma_{0}(t)\big)\in S_{r_{j}}\ \text{ for every }j\in\{0,\ldots,\ell\}\,\big\}.$
By construction $\gamma_{\ell}\in\mathcal{G}_{r_{\ell}}$. Therefore, there exists $g_{\ell}:J_{r_{\ell}}\to\mathbb{R}$ such that $\gamma_{\ell}(t)=(t,g_{\ell}(t))$. Thus, $\gamma_{\ell}(J_{r_{\ell}})=\{(t,g_{\ell}(t));t\in J_{r_{\ell}}\}\subset S_{r_{\ell}}$, and from (22) we obtain that $(\mathcal{T}_{\ell})_{\ell\in\mathbb{N}}$ is a nested sequence of non-empty compact sets, hence
$\bigcap_{\ell=0}^{\infty}\mathcal{T}_{\ell}\neq\emptyset.$
Choose $t_{0}\in\bigcap_{\ell=0}^{\infty}\mathcal{T}_{\ell}$ and set $x:=\gamma_{0}(t_{0})=(t_{0},0)$. Then, for every $\ell\in\mathbb{N}$,
$T_{\omega}^{\ell}(x)=T_{\omega}^{\ell}\big(\gamma_{0}(t_{0})\big)\in S_{r_{\ell}},$
which proves Step 2.
Step 3.
We show that $h_{\mathrm{top}}(\Theta\mid\mathbb{P})>0$, which completes the proof of the proposition.
Given $\omega\in\Omega$, define
$X_{\omega}:=\bigcap_{n=0}^{\infty}T_{\omega}^{-n}\left(J_{1}\sqcup J_{2}\right).$
By Step 2, $X_{\omega}\neq\emptyset$, and the coding map
$h_{\omega}\colon X_{\omega}$ $\rightarrow\{1,2\}^{\mathbb{N}}$
$x$ $\mapsto\big(J(T_{\omega}^{\,i}x)\big)_{i\geq 0},$
where $J(z)=i$ if $z\in J_{i}$ for $i\in\{1,2\}$, is surjective.
Set $X_{\Omega}:=\{(\omega,x);\ \omega\in\Omega\ \text{and}\ x\in X_{\omega}\}.$ Define the forward-time semi-conjugacy
$\pi\colon(\omega,x)\in X_{\Omega}\mapsto(\omega,h_{\omega}(x))\in\Omega\times\{1,2\}^{\mathbb{N}}.$
Let $\sigma\colon\{1,2\}^{\mathbb{N}}\to\{1,2\}^{\mathbb{N}}$ be the left shift and set
$\sigma_{\Omega}\colon(\omega,s)\in\Omega\times\{1,2\}^{\mathbb{N}}\to(A\omega,\sigma(s))\in\Omega\times\{1,2\}^{\mathbb{N}}.$
By construction, $\pi\circ\Theta=\sigma_{\Omega}\circ\pi.$ Since $\pi$ is a surjective factor map and $\Theta$ and $\sigma_{\Omega}$ share the same base $(\Omega,A)$, we obtain
$h_{\mathrm{top}}(\Theta\mid\mathbb{P})\geq h_{\mathrm{top}}(\sigma_{\Omega}\mid\mathbb{P})=h_{\mathrm{top}}(\sigma)=\log 2>0.$
Acknowledgements
MMC acknowledges useful early discussions with Anthony Quas and Yuri Lima. GF and MMC are supported by an Australian Research Council Laureate Fellowship (FL230100088).